### v1

# femtoPro: Real-time linear and nonlinear optics simulations

femtoPro: Real-time linear and nonlinear optics simulations
[…]
journal: josabarticletype: Research Article
[…]
[SI]femtoPro_SI
{abstract*}
Real-time optics and spectroscopy simulations ideally provide results at update rates of  120 Hz or more without any noticeable delay between changing input parameters and the calculated results. Such calculations require models of sufficient speed yet adequate level of detail in the physical approximations to contain the essential features of the simulated phenomena. We discuss a representation of femtosecond laser pulses in which fast phase oscillations due to carrier frequency and due to spatial propagation are separated out and amplitude modulations due to Gaussian beam propagation are also separated and treated explicitly. We derive simplified expressions for the spatial modulations of laser beams. Further, we derive visibility and beam-overlap factors describing multi-pulse interference. We obtain simplified expressions for radius and curvature of nonlinear signal beams in the case of fundamental beams with different convergence, different beam waist, and imperfect mutual overlap. The described model is implemented in the virtual-reality laser laboratory simulation “femtoPro,” but the derived equations can be used independently for other applications.
1 Introduction
Computer simulations of optical phenomena and light–matter interaction are ubiquitous across the scientific disciplines because light fields are elementary probes of nature. In particular in the natural sciences, lasers are used as tools in the realms of frequency-resolved or time-resolved spectroscopy and microscopy applications. The particular optical setups of such techniques may require complex and precise configurations. Planning experiments and analyzing experimental data is hence generally carried out with the help of appropriate computer simulation techniques. For example, optical systems are designed using geometrical ray tracing [1, 2]. If more precision is required, finite-difference methods are employed for numerical approximations of wave optics [3, 4, 5, 6, 7, 8]. Similarly, the simulation of light–matter interactions can be performed at various levels of accuracy and computational cost, including light-induced quantum dynamics [9, 10, 11, 12, 13].
While it is generally desirable to implement procedures and algorithms as efficiently as possible for a given approach, the absolute computation time for nonlinear spectroscopy simulations is typically not the main priority because it is not critical whether results are obtained after a few seconds, minutes, or hours if the desired accuracy is reached in the end. Real-time calculations are not required under such normal-use circumstances.
We have recently introduced an immersive virtual-reality (VR) simulator of an ultrafast laser laboratory (“femtoPro”) [14, 15]. It can be used to provide practical training to students or researchers and allows them to obtain practical expertise in the fundamentals of optics, basic alignment procedures, all the way up to building and using advanced time-resolved spectroscopy setups. For this purpose, real-time simulations are required of linear and nonlinear optical phenomena. In this context, “real time” does not signify the native femtosecond timescale of the photoinduced dynamics, but it means that the total calculation time for a complete spectroscopy experiment should be on the order of $8\,\mathrm{ms}$ on a typical VR headset such that frame update rates of $120\,\mathrm{Hz}$ can be reached. Such rates reduce motion blur [16] and motion sickness [17], improving user comfort. In other words, the goal is that the user does not notice a timing difference between carrying out any linear or nonlinear optical experiment in a real laboratory (with appropriate data acquisition software) or simulating the same setup in femtoPro. While our motivation for the present work thus stems from implementing this VR simulator, the derived equations are much more general and are applicable in other optics simulation scenarios, outside of a VR context, particularly if fast computation is desired.
Thus, the purpose of the present paper is two-fold: 1) to explain the assumptions underlying the femtoPro software such that users may understand its features and limitations, and 2) to introduce and discuss physical models that can be employed for fast simulations on any platform and for various purposes.
The requirement for a lag-free user experience sets a top priority on the speed of the underlying simulations. This is different from the established simulation tools and protocols listed above. Still, the accuracy needs to be sufficient to provide a realistic simulation of the spatial and temporal propagation of ultrafast laser pulses and their interaction with matter. Some requirements are therefore: We desire to take into account Gaussian beam propagation because finite-focus-size effects are relevant to simulate beam overlap alignment procedures; laser beams may be clipped if they hit the edge of an optical element or an iris; femtosecond pulses are modified by dispersion upon propagation through materials and even more so upon resonant interactions; and second-order processes shall be included to simulate frequency conversion, all in real time as defined above.
The underlying simulation concepts are in principle known from the vast existing scientific literature, in particular from textbooks on general optics [18, 19, 20, 21, 22, 23] and ultrafast spectroscopy [24, 25, 26, 27, 28, 29, 30, 31, 32]. However, the real-time simulation requirement necessitates a careful balance between speed and accuracy. In the present work, we derive simplified expressions that facilitate real-time optics simulations.
We start by defining our representation of the electric fields of laser pulses (Section 2). Then we discuss the spatial modulations of laser beams resulting from geometric effects of optical elements and multi-beam interference (Section 3), followed by spatial and spectral–temporal modulations arising from linear and nonlinear light–matter interaction (Section 4) and the graphical representation of laser beams (Section 5). Exemplary simulation results are shown in Section 6 before we conclude in Section 7.
2 Laser pulses
The representation of electric fields of laser pulses is treated in many comprehensive books on general optics [18, 19, 20, 21, 22, 23] and ultrafast time-resolved spectroscopy [24, 25, 26, 27, 28, 29, 30, 31, 32]. Here we seek a representation that is suitable for efficient calculations in computational models. Since we need to describe both spatial and spectral–temporal properties, we have to separate the various aspects in a convenient manner. We define and derive all relevant expressions in great detail in Section LABEL:SISecElectricField of Supplementary Information (SI) and only quote the main results here.
As a result of the derivations, we find that the electric field of a laser pulse at position $\mathbf{r}$ as a function of spatial coordinates $x$, $y$ and $z$ and of time $t$ is given by
$E_{\text{prop}}^{+}(x,y,z,t)=\sqrt{\frac{2}{\pi}}\frac{1}{w(z)}\exp\left[-\frac{(x-x_{0})^{2}+(y-y_{0})^{2}}{w^{2}(z)}\right]e^{i\mathbf{k}\cdot\mathbf{r}}\sqrt{\frac{S}{\delta t}}\tilde{E}_{t}(j)e^{-i\omega_{0}(t-T)}$ (1)
according to Eq. (LABEL:SIEqEPropXYZT) in the SI, where the “$+$” superscript in $E_{\rm{prop}}^{+}$ indicates that this is a complex-valued representation resulting from the positive-frequency part of the full field only and the subscript “prop” indicates that the pulse has propagated through space along the $\hat{\mathbf{z}}$ direction requiring the propagation time $T$. The beam radius at each $z$ position is given by $w(z)$ (with details on the evolution of $w(z)$ as a function of $z$ provided in Section 5) and lateral translations of the beam along $\hat{\mathbf{x}}$ and $\hat{\mathbf{y}}$ by $x_{0}$ and $y_{0}$, respectively. Further, $\mathbf{k}$ denotes the wave vector, $\omega_{0}$ the center frequency, $\delta t$ the sampling step size in time domain, $S$ a scaling factor proportional to pulse energy, and $\tilde{E}_{t}(j)$ the numerical array representation of the complex-valued temporal envelope (including amplitude and phase terms) for individual sampling points that are indexed by $j$.
Analogously, we obtain the temporally propagated spatial–spectral field from Eq. (LABEL:SIEqEPropXYZOmega),
$E_{\textrm{prop}}^{+}(x,y,z,\omega)=\sqrt{\frac{2}{\pi}}\frac{1}{w(z)}\exp\left[-\frac{(x-x_{0})^{2}+(y-y_{0})^{2}}{w^{2}(z)}\right]e^{i\mathbf{k}\cdot\mathbf{r}}\sqrt{\frac{S}{\delta\omega}}\tilde{E}_{\omega}(j)e^{i\omega T},$ (2)
with analogous definitions, and where $\delta\omega$ describes the sampling step size in frequency domain.
3 Geometric effects
3.1 Finite aperture
Apertures lead to the clipping of laser beams. Recently, analytical propagation formulas have been derived for truncated Gaussian beams [33]. Here we ignore diffraction for reasons of computational speed. If we removed the Gaussian beam properties (2) at the same time, that would correspond to a transition from wave optics to the limit of geometrical optics. The latter is not sufficient, however, if we want to describe interference phenomena and allow for the quantitative treatment of spatial beam overlap effects in nonlinear optical phenomena. Thus, in the present work we suggest an intermediate regime where we treat some phenomena in the limit of geometrical optics and some phenomena using wave optics.
Concerning apertures, we select a treatment by geometrical optics. This has the advantage that we do not have to follow various diffracted beamlets that might propagate in various different directions, in particular after interaction with subsequent optical elements. Instead, we can continue to describe the transmitted beam as a single entity with one limited set of parameters. While this is a limitation that removes some phenomena from being treated with the model, it strongly reduces complexity and facilitates real-time simulations.
Despite this approximation, we seek a representation that is as faithful to reality as possible under this approximation. Thus, a laser beam emerging from any open aperture should not have a cross section that extends beyond the hard limits of the aperture. Such an aperture could for example be an iris that is opened or closed by users either to clip laser beams on purpose or to act as alignment tools. The laser beam then would be centered onto at least two subsequent irises, and after the alignment procedure is completed, the irises would be opened to let the full beam pass. In that case, then, the final beam propagation would again not be limited by diffraction for the “real” experiment, justifying the limit of geometrical optics, while the clipping effects will be captured at least qualitatively or semi-quantitatively (to be discussed below) during the simulation of alignment procedures. As a second possibility, finite-aperture effects become relevant when a laser beam is larger than the optical element it hits, or when it hits the element at its edge. This might happen for misaligned mirrors or lenses that are hit by the laser at their edges instead of in the center. While such situations should be simulated at least qualitatively in a real-time optics education scenario, they should typically not arise in a final, correctly set up experiment. Thus, again, it seems justified to employ a simplified geometrical optics treatment.
The approximation of geometrical optics entails that our model cannot be used to simulate diffraction gratings. This is a current limitation (see Section LABEL:SISecFeaturesAndLimitations in the SI for an overview of features and limitations).
In order to ensure that the Gaussian beam that is transmitted through an aperture does not have a cross section that extends beyond the aperture limits, one has to evaluate the geometrical common cross section of the laser beam and the aperture. At normal incidence, one requires the common area of two circles formed by the aperture and the laser cross section, in which one then has to fit a third circle that does not extend beyond the common limits. For nonlinear signal generation from at least two separate beams, however, one would already have to take into account at least three circles, i.e., one for the aperture and two for the incident beams. The problem of the common intersection of three circles is already quite involved with an analytical solution reported only in 2006 [34]. Anyway, the laser beam will, in general, not hit the aperture at normal incidence. In that case, one would have to consider the common overlap of three ellipses that is even more complicated than that of three circles because of the many types of overlap cases that have to be distinguished [35, 36]. Since we ignore astigmatism, however, it is enough to consider a suitably adapted scenario with circular symmetry.
We have adopted the “mixed” solution that evaluates overlaps between several beams, such as in interference evaluation or nonlinear signal generation, on the basis of overlapping Gaussian functions; as a second component of our approach, we then treat the transmission through a hard aperture in a purely geometrical fashion to ensure that “no” laser beam profile extends beyond the open aperture. The term “no” is set in quotation marks here because the visual appearance of the laser beam, when rendered as a three-dimensional object up to its beam radius, will indeed not be larger than the transmission aperture. On the other hand, since we treat the beam as a Gaussian, this still allows its radial distribution to extend further outwards than the aperture opening. This is a consequence of the hybrid approach, adopted in this work, between geometrical optics and wave optics when diffraction is ignored. Thus, we take the beam radius as the geometric limit. When we then choose to visualize laser beams via plotting their beam radius, an intuitively consistent behavior emerges if the beam waist is taken to represent the “edge.”
For treating the transmission through optics, we have to consider finite incidence angles. For example, mirrors are routinely employed at $\approx 45^{\circ}$ incidence angle to deflect beams, and anyway we cannot avoid that the beam hits a general optical element (GOE) at an arbitrary angle, so the model needs to deal with all situations. Thus we may not assume the “paraxial” limit, and we employ vector calculus for an analysis independent of any particular chosen coordinate system.
The relevant geometry is illustrated in Fig. 1. The incident laser beam (from the left, potentially arriving at an additional angle elevated out of the drawing plane) is given by a point of origin in world coordinates, $\mathbf{r}_{\text{in}}$, in our evaluation logic usually its starting point at the previous GOE; a unit directional vector $\hat{\mathbf{k}}_{\text{in}}=\mathbf{k}_{\text{in}}/\left\lvert\mathbf{k}_{\text{in}}\right\rvert$, where $\mathbf{k}_{\text{in}}$ is the incident wave vector; and a beam radius $w_{\text{in}}$ at the point of intersection with the current GOE. The current GOE is defined by its aperture center in world coordinates, $\mathbf{o}$, a normal vector describing the orientation of the “active” aperture plane, $\hat{\mathbf{n}}$, and an aperture radius $a$.
Figure 1: Aperture geometry. The input beam (red) propagates along a central “ray” (solid) with lateral width determined by its beam radius (dashed–dotted) until it intersects at an incidence angle $\alpha$ with the central active plane of a general optical element (GOE) that is assumed, without loss of generality, to be normal to the “optical axis” (dashed black). The output beam (green) is likewise defined by a central ray (solid) and radius (dashed–dotted). All relevant scalar and vectorial quantities are marked in blue as defined in the text.
We first determine the propagation vector, $\mathbf{L}$, that extends from the origin of the laser beam at the previous GOE to the intersection with the active plane of the current GOE [37],
$\mathbf{L}=\frac{(\mathbf{o}-\mathbf{r}_{\text{in}})\cdot\hat{\mathbf{n}}}{\hat{\mathbf{k}}_{\text{in}}\cdot\hat{\mathbf{n}}}\,\hat{\mathbf{k}}_{\text{in}}\quad\text{if }\hat{\mathbf{k}}_{\text{in}}\cdot\hat{\mathbf{n}}\neq 0.$ (3)
If $\hat{\mathbf{k}}_{\text{in}}\cdot\hat{\mathbf{n}}=0$, the laser beam propagates parallel to the GOE’s active plane, and we have to distinguish further cases. If, in that case, the numerator in Eq. (3) is not equal to zero, the laser beam is parallel to but outside of the surface, so that there is no intersection. If the numerator and the denominator in Eq. (3) are equal to zero, then the laser beam propagates within the surface, i.e., the GOE is hit exactly from the side. We then define the propagation vector such that it propagates the laser beam from its origin to the location closest to the center of the GOE,
$\mathbf{L}=\mathbf{o}-\mathbf{r}_{\text{in}}-\text{proj}_{\hat{\mathbf{k}}_{\text{in}}}(\mathbf{o}-\mathbf{r}_{\text{in}}),$ (4)
with the vector projection
$\text{proj}_{\mathbf{b}}\mathbf{a}=\left(\mathbf{a}\cdot\frac{\mathbf{b}}{\left\lvert\mathbf{b}\right\rvert}\right)\frac{\mathbf{b}}{\left\lvert\mathbf{b}\right\rvert}$ (5)
of vector $\mathbf{a}$ onto vector $\mathbf{b}$.
The propagation length $L=\left\lvert\mathbf{L}\right\rvert$ enters Eq. (LABEL:SIEqPropagationTime) in the SI for the calculation of the propagation time $T$. From the previous origin of the laser and the propagation vector we obtain the “hit” point of intersection with the current GOE, in world coordinates, as
$\mathbf{h}=\mathbf{r}_{\text{in}}+\mathbf{L},$ (6)
and as
$\mathbf{d}=\mathbf{h}-\mathbf{o}$ (7)
in a GOE-local coordinate system with the origin at the center of the GOE, where we further define the scalar intersection distance $d=\left\lvert\mathbf{d}\right\rvert$ from the GOE center. The vectorial quantity $\mathbf{d}$ enters the calculation of directional change by a GOE with focal length not equal to zero, as will be derived below. In the present context, we project the incident beam radius, $w_{\text{in}}$, along direction $\mathbf{d}$ to obtain the projected beam radius, $b_{\text{in}}$. For this, we observe in Fig. 1 that $\cos\alpha=w_{\text{in}}/b_{\text{in}}$ but also that $\cos\alpha=\hat{\mathbf{n}}\cdot\hat{\mathbf{k}}_{\text{in}}$. Eliminating $\alpha$ provides
$b_{\text{in}}=\frac{w_{\text{in}}}{\left\lvert\hat{\mathbf{n}}\cdot\hat{\mathbf{k}}_{\text{in}}\right\rvert},$ (8)
where we introduced the absolute magnitude operator such that it works also if the surface normal $\hat{\mathbf{n}}$ is inverted. In the case of $\hat{\mathbf{n}}\cdot\hat{\mathbf{k}}_{\text{in}}=0$, as noted above, the beam travels within the plane, and the projected radius is infinite as it should be. This leads to the further treatment as case 2 in the list below.
Now we have obtained the three quantities, aperture radius $a$, intersection distance $d$, and projected beam radius $b_{\text{in}}$, that are required for the following case distinction based on their quantitative relation:
  1. 1.
    $a\geq d+b_{\text{in}}$: The laser beam is completely within the aperture;
  2. 2.
    $b_{\text{in}}>d+a$: The laser beam overfills completely the aperture;
  3. 3.
    $d>b_{\text{in}}+a$: The laser beam is completely outside the aperture;
  4. 4.
    Else: The laser beam is partially within the aperture.
Let us evaluate, for each case, the projected radius of the new beam after the aperture, $b_{\text{out}}$, along direction $\mathbf{d}$; the origin of the new beam after the aperture in world coordinates, $\mathbf{r}_{\text{out}}$; and the direction of the new beam after the aperture, $\hat{\mathbf{k}}_{\text{out}}$. Finally, we obtain the new beam radius perpendicular to the new direction of propagation, $w_{\text{out}}$, that is given in all cases, analogous to Eq. (8), by
$w_{\text{out}}=b_{\text{out}}\left\lvert\hat{\mathbf{n}}\cdot\hat{\mathbf{k}}_{\text{out}}\right\rvert.$ (9)
Note that in such a treatment, we consider the beam radius as projected along direction $\mathbf{d}$ only, and we ignore the shape and extension in the orthogonal direction. This is an approximation that works well for meridional rays but less so for skew rays. Nevertheless, we follow this approach because of our choice to ignore astigmatism for computational efficiency [14], and thus one may define just one particular projection coordinate (not two) to represent a beam radius and shape. Using the meridional-ray approximation leads to the desired effect that no transmitted beam is larger than any circular open aperture.
Case 1 ($a\geq d+b_{\text{in}}$) in the list above corresponds to the “normal” alignment goal of having an unclipped beam with an open iris aperture (or a fully reflected beam off a mirror). Then the starting point of the new beam is equal to the intersection point of the incident beam,
$\mathbf{r}_{\text{out}}=\mathbf{h},$ (10)
the new beam direction is given by [38]
$\hat{\mathbf{k}}_{\text{out}}=\frac{\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{d}}{f}}{\left\lvert\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{d}}{f}\right\rvert},$ (11)
and the projected radius is simply
$b_{\text{out}}=b_{\text{in}}.$ (12)
In case 2 ($b_{\text{in}}>d+a$), the clipped beam passes, symmetrically distributed, exactly through the center of the GOE, independently of the local intersection coordinate $\mathbf{d}$ that was calculated for the center of the unclipped beam. Note that in this geometric treatment, as motivated at the beginning of this section, we do not take into account the Gaussian variation of the intensity within the lateral cross section, and thus we arrive at the statement of the symmetric distribution around the GOE center $\mathbf{o}$. Thus, this case corresponds to a “center ray” in geometrical optics that is unaffected by any focal length, and hence we obtain
$\mathbf{r}_{\text{out}}=\mathbf{o}$ (13)
for the origin,
$\hat{\mathbf{k}}_{\text{out}}=\hat{\mathbf{k}}_{\text{in}}$ (14)
for the direction, and
$b_{\text{out}}=a$ (15)
for the projected radius of the new beam.
Case 3 ($d>b_{\text{in}}+a$) is simple to treat because then no beam is transferred.
Finally, in case 4 (“else”), the beam is clipped partially. From Fig. 1, one finds $2b_{\text{out}}=a-d+b_{\text{in}}$ such that the projected radius is
$b_{\text{out}}=\frac{1}{2}(a-d+b_{\text{in}}).$ (16)
If the beam shall indeed arise from the common overlap between original projected radius and aperture radius, we have to laterally “move” the new beam’s origin such that it is in the center of the new beam. From Fig. 1, one obtains $p=\left\lvert\mathbf{p}\right\rvert=d-b_{\text{in}}+b_{\text{out}}=\frac{1}{2}(a+d-b_{\text{in}})$ and, as a GOE-local vector,
$\mathbf{p}=p\frac{\mathbf{d}}{d}=\frac{1}{2}(d+a-b_{\text{in}})\frac{\mathbf{d}}{d}$ (17)
such that
$\mathbf{r}_{\text{out}}=\mathbf{o}+\mathbf{p}$ (18)
and
$\hat{\mathbf{k}}_{\text{out}}=\frac{\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{p}}{f}}{\left\lvert\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{p}}{f}\right\rvert}.$ (19)
Note that now the effective “distance” entering in Eq. (19) is $\mathbf{p}$ rather than $\mathbf{d}$ as in Eq. (11). We illustrate the case of partial beam clipping in Fig. 2, where we vary the lateral shift of a beam incident on a GOE with a finite aperture while displaying the waist and the angle of an output beam. The radii of the input beam and the aperture are set to 3 mm. To invoke a change of the angle of the output beam, we set the focal length of the GOE to $f=37.5$ cm. We display two subcases. In case of a small angle of incidence (5.71), the output beam waist decreases linearly when shifting the input beam away from the center of the aperture [Fig. 2(a)]. The output beam angle also changes linearly. In case of a steeper angle of incidence (45) as shown in Fig. 2(b), the radius and angle of the outgoing beam remain constant in a region around the center of the aperture. When the incident beam is further laterally shifted, clipping again results in a change of radius and angle of the outgoing beam.
Figure 2: Output beam from a finite aperture. The lateral shift of an input beam relative to the optical axis is varied under a given incidence angle. The beam waist of the input beam and the aperture radius are both chosen to be 3 mm while the focal length of the GOE is set to 37.5 cm. The beam waist and the angle of the output beam is shown for an incidence angle of (a) 5.71 and (b) 45. The incidence angles are indicated as horizontal dashed lines.
Finally, we need to consider the reduction in pulse energy due to beam clipping caused by the finite aperture. Above, we made the approximation of uniform spatial profiles for generating the geometric parameters of the new beam. Let us first consider what would happen at the same level of approximation for the transmitted pulse energy (which, as we will see, is not what we implemented). In this simplest approximation of a “flat-top” beam, we could relate the output energy, $W_{\text{out}}$, to the incident pulse energy, $W_{\text{in}}$, by considering the ratio of the projected cross-sectional areas at the aperture,
$W_{\text{out}}=\frac{\pi b_{\text{out}}^{2}}{\pi b_{\text{in}}^{2}}W_{\text{in}}.$ (20)
Note that one should not use the orthogonal beam radii $w_{\text{in}}$ and $w_{\text{out}}$ in this context because they are different already if the output beam direction $\hat{\mathbf{k}}_{\text{out}}$ is different from the incidence beam’s direction $\hat{\mathbf{k}}_{\text{in}}$ due to an off-axis focusing element with focal length $f\neq 0$, i.e., the pulse energy does not change but only the intensity changes due to the changed area.
While Eq. (20) describes the situation for a uniform spatial intensity profile, this does not correspond to the properties of a Gaussian beam in which the intensity is higher in the center than at the edge and which is what we introduced in Eq. (1) and Eq. (2). The discrepancy would be noticeable and is relevant if the simulation model is used, for example, for didactic training in the context of beam alignment onto an aperture. In those cases, the goal is to align the position of a laser beam with the center of a given aperture (or vice versa, center an aperture on a given laser beam) by monitoring the energy changes detected with a power meter after the aperture when moving the beam (or aperture) laterally. Assume now that the laser overfills the aperture (case 2 in the list above). In that case, the transmitted cross-sectional area would not change, and neither would the transmitted pulse energy according to simplistic Eq. (20). Thus, we seek a more accurate calculation of the transmitted energy.
More accurately, we might alternatively consider, instead of Eq. (20) from above, equation (14) from reference [39] that derives, for normal incidence at which $b_{\text{in}}=w$, the following relation:
$W_{\text{out}}=W_{\text{in}}\exp\left(-\frac{2d^{2}}{w^{2}}\right)\times\sum_{k=0}^{\infty}\left\{\frac{2^{k}d^{2k}}{w^{2k}k!}\left[1-\exp\left(-\frac{2a^{2}}{w^{2}}\right)\sum_{i=0}^{k}\frac{2^{i}a^{2i}}{w^{2i}i!}\right]\right\}.$ (21)
It is discussed in reference [39] that for central beams, i.e., $d=0$, one obtains the more well-known relation
$W_{\text{out}}(d=0)=W_{\text{in}}\left[1-\exp\left(-\frac{2a^{2}}{w^{2}}\right)\right],$ (22)
but this, on its own, is not a useful limit for us either because we want to deal with the particular situation of laterally displaced beams.
Thus we evaluate the transmitted energy numerically using precalculated lookup tables as derived in Section LABEL:SISecGaussianTransmissionThroughCircularAperture in the SI. This allows us to assign transmission values fast for any given combination of beam radius $w$, aperture radius $a$, and displacement $d$. Note that we do not change the pulse energy directly, but rather the pulse-energy scaling factor $S$ from Eq. (LABEL:SIEqEnergyIsScaleTimesNormSquared). The scaling factor as a function of the aperture radius is displayed in Fig. 3(a), where the behavior of Gaussian beams is evident by an $S$ less than one in the case of $w=a$. In Fig. 3(b), we laterally shift the input beam relative to an aperture with size $a=3w$. Note that we capture the beams geometrically up to their beam radius, i.e., we treat Gaussian beams, which theoretically go to infinity, as hard-cut structures. This means that if the beam is shifted away from the edge of the aperture by a distance larger than $w$, it is treated as no longer hitting the aperture, leading to zero energy which in turn results in the small “step” visible in Fig. 3(b) at a lateral shift of $4w$. This provides a quantitative estimate for the error at the chosen level of approximation of our model.
Figure 3: Pulse-energy scaling factor S of a Gaussian beam incident on an aperture. (a) Scaling factor as a function of the aperture size. (b) Scaling factor as a function of the lateral shift relative to an aperture with a size of $3w$.
3.2 N-beam interference
We now analyze the interference of $N$ laser pulses, $E_{k,\text{prop}}^{+}(\mathbf{r},\omega)$, $k=\{1,\ldots,N\}$, each given by Eq. (2), on a plane such as a detector. For this purpose, we evaluate the spectral power $P_{\omega}$, as detected by a spectrometer, by integrating the absolute magnitude squared of the total field over transverse coordinates [14],
$P_{\omega}(j)=$ $\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\left\lvert\sum_{k=1}^{N}E_{k,\text{prop}}^{+}(\mathbf{r},\omega)\right\rvert^{2}\,\mathrm{d}x\,\mathrm{d}y$ (23)
$=$ $\sum_{k=1}^{N}\frac{S_{k}}{\delta\omega}\left\lvert\tilde{E}_{k,\omega}(j)\right\rvert^{2}+\sum_{k=2}^{N}\sum_{l=1}^{k-1}2\operatorname{Re}\left\{\eta_{k,l}\frac{\sqrt{S_{k}S_{l}}}{\delta\omega}\tilde{E}_{k,\omega}(j)\tilde{E}_{l,\omega}^{*}(j)e^{i\omega(T_{k}-T_{l})}\right\}.$ (24)
Here we derive the factor $\eta_{k,l}$ regulating interference visibility between two fields with indices $k$ and $l$. Note that we do not want to describe the spatially resolved visibility of interference fringes, but their integrated effect when evaluating a finite detector area. Then, interference is noticeable if the spatial fringe spacing is large enough such that it does not average out when integrating over all fringes. Considering the field definition in Eq. (2), we have to evaluate
$\begin{split}\eta_{k,l}=\,&\frac{2}{\pi w_{k}w_{l}}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\exp\left[-\frac{(x-x_{k})^{2}+(y-y_{k})^{2}}{w_{k}^{2}}\right]\exp\left[-\frac{(x-x_{l})^{2}+(y-y_{l})^{2}}{w_{l}^{2}}\right]\\ &\times e^{i\Delta\mathbf{k}\cdot\mathbf{r}}\,\mathrm{d}x\,\mathrm{d}y\end{split}$ (25)
with the wave-vector mismatch
$\Delta\mathbf{k}=\mathbf{k}_{k}-\mathbf{k}_{l}.$ (26)
We proceed in the coordinate system in which the $\hat{\mathbf{z}}$ axis is parallel to the average of the incident wave vectors. As will be derived below, interference is visible only for small phase mismatch, and thus the approximation is valid that the beam positions, $(x_{k},y_{k})$ and $(x_{l},y_{l})$, and radii, $w_{k}$ and $w_{l}$, need not be transformed because the average direction is almost identical to either of the incident directions. The actual incidence angle onto the detector plane is not decisive in first order because the underlying spatial interference pattern is retained, and thus integration over the tilted pattern provides a comparable total power as integration over the normal-incidence pattern.
In Section LABEL:SISecOverlapOfTwoDimensionalGaussianFunctions of the SI, we show that the product of two two-dimensional Gaussian cross-sections is another Gaussian cross-section, and we derive its resulting product width parameters $\alpha_{\text{p}}$ and $\beta_{\text{p}}$, transverse position $(x_{\text{p}},y_{\text{p}})$, and product amplitude $A_{\text{p}}$ (see SI for definitions). Using those results for unity incident amplitudes, we have to evaluate
$\eta_{k,l}=\frac{2A_{\text{p}}}{\pi w_{k}w_{l}}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\exp\left[-\alpha_{\text{p}}(x-x_{\text{p}})^{2}-\beta_{\text{p}}(y-y_{\text{p}})^{2}\right]e^{i(\Delta k_{x}x+\Delta k_{y}y)}\,\mathrm{d}x\,\mathrm{d}y.$ (27)
This integral is solved in Section LABEL:SISecInterferenceVisibilityFactor of the SI, leading to the already reported result [14]
$\begin{split}\eta_{k,l}=&\frac{2w_{k}w_{l}}{w_{k}^{2}+w_{l}^{2}}\exp\left[-\frac{(x_{k}-x_{l})^{2}+(y_{k}-y_{l})^{2}}{w_{k}^{2}+w_{l}^{2}}\right]\\ &\times e^{i(\Delta k_{x}x_{\text{p}}+\Delta k_{y}y_{\text{p}})}\exp\left[-\frac{\Delta k_{x}^{2}+\Delta k_{y}^{2}}{2\left(\frac{1}{w_{k}^{2}}+\frac{1}{w_{l}^{2}}\right)}\right].\end{split}$ (28)
Note that interference visibility in the literature is often treated in terms of the coherence functions of the light fields. This is taken into account in Eq. (24) by the specific properties of the spectral envelope functions. We discuss a frequency-resolved version of interference in which coherence pertains for arbitrarily long differences in propagation distance (by definition of a spectral field). For the spectrally integrated result, one then obtains the usual limit of optical coherence and interference visibility that disappears when fields are separated longitudinally by a distance larger than their coherence length (or, for bandwidth-limited ultrashort pulses, by delay times larger than their pulse duration).
In Fig. 4(a), we illustrate the dependence of the interference visibility $\eta_{1,2}$ between two beams on various parameters. In the first example [Fig. 4(a)], one beam remains constant and the second beam is shifted laterally with otherwise identical parameters. We show three exemplary curves for the cases in which both beams have a radius of 1 mm (blue), 2 mm (red), and 4 mm (green). It is evident that the interference visibility can be maintained over a larger lateral shift range in case of larger beam radii. In Fig. 4(b), we show the influence of the wave-vector mismatch on the interference visibility by varying the propagation angle of the second beam relative to the first. Here, the angular mismatch has less impact on the interference visibility when both beams have smaller radii, which reflects the general behavior of the last exponential function in Eq. (28).
Figure 4: Interference visibility factor $\eta_{1,2}$ between two light fields. (a) Interference visibility resulting from the second beam being laterally shifted relative to the first beam where both beams have a radius of 1 mm (blue), 2 mm (red) or 4 mm (green). (b) Dependence of the interference visibility on the relative incidence angle of the second beam [colors correspond to the same beam radii as in (a)]. (c) Interference visibility of two beams incident on an aperture with varying radius. Here, both beams have a radius of 4 mm and one of the beams is laterally shifted by 1 mm from the center of the aperture at normal incidence while the other beam is laterally shifted by 4 mm relative to the first beam and has an incidence angle of 0.01 (fixed parameters).
In Section LABEL:SISecInterferenceVisibilityFactor of the SI, we discuss quantitatively how interference visibility can be increased by closing an entrance aperture $a$ directly in front of a spectrometer, and we find the conditions for maximal angular mismatch and transverse misalignment between incident beams that still ensure interference. In Fig. 4(c), we display the interference visibility factor in the case where two beams with a radius of 4 mm are incident on an aperture while the size of the aperture is changed. One of the beams is laterally shifted by 1 mm relative to the center of the aperture at normal incidence whereas the other beam has an exemplary lateral shift of 4 mm relative to the first beam and is incident at an angle of $0.01^{\circ}$. The two beams hit the aperture at a distance of 1 m from their origin, resulting in the second beam being displaced in the plane of the iris by 4.8 mm relative to the center of the iris. This configuration leads to a beam overlap region that is displaced relative to the center of the iris, as indicated in the inset of Fig. 4(c). This example shows that closing the aperture (initially considered to have an infinite opening radius) leads to an increase of the interferometric contrast as those parts of the beams that do not overlap get truncated. However, when closing the aperture further, the overlapping parts get truncated as well, which leads to reduction in contrast such that there is an optimum aperture opening. Note that the slight kink at an aperture radius of 3 mm is due to the transition in how the beam cutting is calculated as discussed in connection with Fig. 3(b).
4 Modulations from light–matter interaction
4.1 First-order non-resonant response
The electric-field modifications by linear response of matter is most conveniently implemented by multiplying the incident frequency-domain field, $E_{\text{in}}(\omega)$, with a frequency-domain linear modulation function, $M(\omega)$, to get the output field,
$E_{\text{out}}(\omega)=M(\omega)E_{\text{in}}(\omega),$ (29)
after transmission through the material [25]. In the algorithmic realization, we use, equivalently, the $\tilde{E}_{\omega}(j)$ array. For treating non-resonant dispersion in matter, we follow the standard convention to perform a Taylor expansion of the dispersive spectral phase $\Phi_{\text{disp}}(\omega)$ in Section LABEL:SISecTaylorExpansionOfDispersivePhase of the SI, resulting in Taylor coefficients $b_{j,\text{disp}}$ of $j$th order.
Considering the effect on $\Phi_{\text{disp}}(\omega)$ of the zeroth- and first-order Taylor coefficients explicitly and rearranging, we get
$b_{0}+b_{1}(\omega-\omega_{0})=\frac{L\omega_{0}(n_{0}-n_{\text{gr}})}{c}+\frac{Ln_{\text{gr}}}{c}\omega$ (30)
for propagation through a medium of length $L$ with refractive index $n_{0}$ and group index $n_{\text{gr}}$, wherein the first term is a frequency-uniform phase and the second term is linear in $\omega$. Note that we treat the propagation time $T$, arising from the geometry of free-space propagation, separately according to Eq. (LABEL:SIEqEOmegaProp) from the SI. Thus, when transmitting through a material of thickness $L$, we have to replace $T$ with
$T_{\text{total}}=T+\frac{L}{c}(n_{\text{gr}}-1)$ (31)
and apply only the remaining terms $b_{2,\text{disp}}$ and $b_{3,\text{disp}}$ in the non-resonant part of the linear modulation function, $M_{\text{non-resonant}}(\omega)$. Note that concerning the geometry, we work in the limit of infinitesimally thin GOEs with one central “active plane,” so that the incident and outgoing beams have already been assumed to travel in vacuum up to and starting from that plane. Thus, we need to subtract $L/c$ from the (modified) propagation time because it has already been accounted for, leading to the term “$-1$” in Eq. (31). In the case of reflection, finite values for the group-delay dispersion or the third-order phase coefficient may be provided directly to describe, e.g., chirped mirrors.
4.2 First-order resonant response
It is convenient to start with a response function in time domain, rather than frequency domain, to treat a resonantly excited material such as an ensemble of molecules in a solvent. The frequency-domain modulation function is then obtained by Fourier transformation via Eq. (LABEL:SIEqEomegaJFourier) and appropriate further scaling, as described below.
The linear temporal response function in a Franck–Condon model under the rotating-wave approximation (RWA) is given by [24]
$S_{\text{RWA}}^{(1)}(t)=\theta(t)\frac{i}{\hbar}\mu_{\text{eg}}^{2}\exp\left(-\gamma t-\frac{1}{2}\Delta\omega^{2}t^{2}\right)e^{-i(\omega_{\text{eg}}-\omega_{0})t}e^{-S_{\text{HR}}}\sum_{j=0}^{N_{\text{vib}}-1}\frac{S_{\text{HR}}^{j}}{j!}e^{-ij\omega_{\text{vib}}t},$ (32)
[…]
We now have to introduce an appropriate scaling of $\mu_{\text{eg}}$ to convert between “physical” quantities and computationally stored arrays. This problem is similar to the scaling of electric fields that we solved by normalizing with respect to a given temporal power and, ultimately, a given pulse energy, via Eq. (LABEL:SIEqTemporalPowerDefinition). Thus we apply a similar strategy here. Furthermore, we seek a behavior of the modulation function such that longer paths through the material lead to stronger absorption fulfilling the Lambert–Beer law for the pulse energy,
$W_{\text{out}}=W_{\text{in}}e^{-\alpha L},$ (33)
with a given extinction coefficient $\alpha$ of the material.
As a first step, we evaluate Eq. (32) on the same time grid as for regular laser pulses [Eq. (LABEL:SIEqt(j))] and set $\mu_{\text{eg}}^{2}/\hbar=1$. The Fourier transform of the result is the (unscaled) frequency-domain linear susceptibility,
$\chi_{\text{unscaled}}(\omega)=\operatorname{\mathfrak{F}}\left[\left.S_{\text{RWA}}^{(1)}(t)\right|_{\mu_{\text{eg}}^{2}/\hbar=1}\right],$ (34)
from which we get the (unscaled) complex index of refraction,
$\tilde{n}_{\text{unscaled}}(\omega)=\sqrt{\chi_{\text{unscaled}}(\omega)+1},$ (35)
where it is understood that the square-root symbol indicates a complex-valued square-root operation. Now we obtain a suitable scaling factor,
$s=\frac{\alpha c}{2\omega_{\alpha}\operatorname{Im}\tilde{n}_{\text{unscaled}}(\omega_{\alpha})},$ (36)
where $\alpha$ is the absorption coefficient of Eq. (33) to which we want to scale the material response and $\omega_{\alpha}$ is the angular frequency at which this absorption coefficient shall be reached (typically a peak of the resulting absorption spectrum that can be found by locating the frequency at which $\operatorname{Im}\tilde{n}_{\text{unscaled}}(\omega)$ has a maximum). Using Eq. (36) allows us to calculate a scaled susceptibility,
$\chi(\omega)=s\chi_{\text{unscaled}}(\omega),$ (37)
and a scaled complex index of refraction,
$\tilde{n}(\omega)=\sqrt{\chi(\omega)+1},$ (38)
and finally the desired modulation function,
$M_{\text{resonant}}(\omega)=\exp\left(i\frac{\tilde{n}(\omega)\omega L}{c}\right),$ (39)
for a material of thickness $L$.
[…]
Figure 5: Linear resonant response. Exemplary modulation function $M_{\text{resonant}}(\omega)$ for an absorbing molecular response (red: real part, blue: imaginary part).
4.3 Second-order response
We describe the non-resonant second-order nonlinear generation of a signal field, $E_{\text{s}}(t)$, in the approximation of being proportional to the square of the sum of incident fields, $E(t)$. Considering two incident fields $E_{1}(t)$ and $E_{2}(t)$ with $E(t)=E_{1}(t)+E_{2}(t)$ and complex envelopes $\tilde{E}_{1}(t)$ and $\tilde{E}_{2}(t)$, this leads to the explicit terms
$\begin{split}E^{2}(t)&=\tilde{E}_{1}^{2}(t)e^{i(2\mathbf{k}_{1}\cdot\mathbf{r}-2\omega_{1}t)}+\text{c.c.}+\tilde{E}_{2}^{2}(t)e^{i(2\mathbf{k}_{2}\cdot\mathbf{r}-2\omega_{2}t)}+\text{c.c.}\\ &+2\tilde{E}_{1}(t)\tilde{E}_{2}(t)e^{i[(\mathbf{k}_{1}+\mathbf{k}_{2})\cdot\mathbf{r}-(\omega_{1}+\omega_{2})t]}+\text{c.c.}\\ &+2\tilde{E}_{1}(t)\tilde{E}_{2}^{*}(t)e^{i[(\mathbf{k}_{1}-\mathbf{k}_{2})\cdot\mathbf{r}-(\omega_{1}-\omega_{2})t]}+\text{c.c.}\\ &+2\left\lvert\tilde{E}_{1}(t)\right\rvert^{2}+2\left\lvert\tilde{E}_{2}(t)\right\rvert^{2},\end{split}$ (40)
where “c.c.” indicates the complex conjugate of the previous term. This is the level of treatment often found in didactic textbooks on second-order response. However, we need to consider also the spatial properties of the beams if we want to simulate and understand the effect of beam alignment on nonlinear phenomena. Here we derive the explicit equations in terms of the computationally stored complex envelope arrays $\tilde{E}_{t}(j)$ and associated parameters, we take into account the spatial beam profile, and we determine the missing proportionality factor between $E_{\text{s}}(t)$ and $E^{2}(t)$.
Equation (40) contains multiplicative electric-field terms. Hence, we use Eq. (1) and multiply the propagated fields,
$\begin{split}&mE_{1,\text{prop}}^{+}(x,y,z,t)E_{2,\text{prop}}^{+}(x,y,z,t)\\ =\,&m\sqrt{\frac{2}{\pi}}\frac{1}{w_{1}(z)}\exp\left[-\frac{(x-x_{1})^{2}+(y-y_{1})^{2}}{w_{1}^{2}(z)}\right]e^{i\mathbf{k}_{1}\cdot\mathbf{r}}\sqrt{\frac{S_{1}}{\delta t}}\tilde{E}_{1,t}(j_{1})e^{-i\omega_{1}(t-T_{1})}\\ &\times\sqrt{\frac{2}{\pi}}\frac{1}{w_{2}(z)}\exp\left[-\frac{(x-x_{2})^{2}+(y-y_{2})^{2}}{w_{2}^{2}(z)}\right]e^{i\mathbf{k}_{2}\cdot\mathbf{r}}\sqrt{\frac{S_{2}}{\delta t}}\tilde{E}_{2,t}(j_{2})e^{-i\omega_{2}(t-T_{2})},\end{split}$ (41)
wherein the multiplication of the two Gaussian exponentials generates a new Gaussian $A_{\text{p}}(x,y)$ according to Eq. (LABEL:SIEqGaussianProduct) of the SI with the product amplitude $A_{\text{p}}$ in Eq. (LABEL:SIEqProductAmplitude) and $A_{1}=A_{2}=1$. Because we consider the stigmatic approximation, we have $w_{1,x}=w_{1,y}$ and $w_{2,x}=w_{2,y}$ according to Eq. (LABEL:SIEqGaussianProduct) to Eq. (LABEL:SIEqProductAmplitude) of the SI. Further, $m=1$ for second-harmonic generation (SHG) but $m=2$ for sum-frequency generation (SFG) according to Eq. (40) due to the two possible permutations $E_{1}E_{2}$ and $E_{2}E_{1}$ that both deliver an SFG field in the same direction.
We want to write the result of Eq. (41) as a new (but not yet correctly energy-scaled) SFG field,
$\begin{split}E_{\text{SFG,prop}}^{+}(x,y,z,t)=\,&\sqrt{\frac{2}{\pi}}\frac{1}{w_{\text{SFG}}}\exp\left[-\frac{(x-x_{\text{SFG}})^{2}+(y-y_{\text{SFG}})^{2}}{w_{\text{SFG}}^{2}}\right]e^{i\mathbf{k}_{\text{SFG}}\cdot\mathbf{r}}\\ &\times\sqrt{\frac{S_{\text{SFG,unscaled}}}{\delta t}}\tilde{E}_{\text{SFG},t}(j_{\text{SFG}})e^{-i\omega_{\text{SFG}}(t-T_{\text{SFG}})},\end{split}$ (42)
with appropriate parameters. Comparing the individual factors of Eq. (41) and Eq. (42), we find:
[…]
$\omega_{\text{SFG}}$ $=\omega_{1}+\omega_{2},$ (46)
$\mathbf{k}_{\text{SFG}}$ $=\mathbf{k}_{1}+\mathbf{k}_{2},$ (47)
$T_{\text{SFG}}$ $=\frac{\omega_{1}T_{1}+\omega_{2}T_{2}}{\omega_{1}+\omega_{2}},$ (48)
$\tilde{E}_{\text{SFG},t}(j_{\text{SFG}})$ $=\tilde{E}_{1,t}(j_{1})\tilde{E}_{2,t}(j_{2}),$ (49)
$\sqrt{\frac{2}{\pi}}\frac{1}{w_{\text{SFG}}}\sqrt{\frac{S_{\text{SFG,unscaled}}}{\delta t}}$ $=m\frac{2}{\pi}\frac{1}{w_{1}(z)w_{2}(z)}A_{\text{p}}\frac{\sqrt{S_{1}S_{2}}}{\delta t}.$ (50)
Note that the beam radius of the SFG beam is smaller than those of the fundamental beams, as expected due to the field multiplication. If both incident beams have the same radius, they are larger than the SFG radius by a factor of $\sqrt{2}$ according to Eq. (45). Solving Eq. (50) for $S_{\text{SFG,unscaled}}$, using the explicit results for $w_{\text{SFG}}$ from Eq. (45) and for $A_{\text{p}}$ from Eq. (LABEL:SIEqProductAmplitude) (with $A_{1}=A_{2}=1$) and simplifying leads to
$S_{\text{SFG,unscaled}}=\frac{2m^{2}}{\pi}\frac{1}{w_{1}^{2}(z)+w_{2}^{2}(z)}\frac{S_{1}S_{2}}{\delta t}\exp\left[-2\frac{(x_{1}-x_{2})^{2}+(y_{1}-y_{2})^{2}}{w_{1}^{2}(z)+w_{2}^{2}(z)}\right].$ (51)
We illustrate how $S_{\text{SFG,unscaled}}$ varies when displacing one of the fundamental beams relative to the other in Fig. 6. Each of the fundamental beams has a pulse energy of 1 mJ and the beam radius of one beam was set to 1 mm (blue), 2 mm (red), and 4 mm (green) while that of the other beam is kept at 1 mm. It is evident that the scaling factor generally shows the largest change with increasing overlap when both beams have small radii. In other words, the SFG intensity at maximum overlap increases with increasing energy density of the fundamental beams.
Figure 6: Beam energy scaling factor for sum-frequency generation. The plots show the dependence of the beam energy scaling factor for sum-frequency generation, $S_{\text{SFG,unscaled}}$, on the lateral shift between two fundamental beams. The beam radius of the first beam was varied (blue: 1 mm, red: 2 mm, green: 4 mm), whereas that of the second beam was held constant (1 mm).
Concerning the multiplication of envelopes in Eq. (49), we have to use the appropriate array indices. Note that we have to add the respective propagation times $T_{1}$ and $T_{2}$ to Eq. (LABEL:SIEqt(j)) before extracting the relevant indices $j_{1}$ and $j_{2}$ of the array elements to be multiplied in Eq. (49). In the multiplication, it will thus happen that a theoretically requested array index is out of range of the available indices $j=\{0,1,\ldots,N_{\text{s}}-1\}$, i.e., the time shift between the two pulses is such that for a time coordinate $t(j_{1})$ at index $j_{1}$ of the first field no “matching” index $j_{2}$ is found because the particular $t(j_{1})$ is not contained in the list for $t(j_{2})$ of the second field. Then we assume the multiplier field to be zero, and thus also the SHG field will vanish at this time coordinate.
Now we have to find an appropriate scaling factor for the SFG field. In principle, proper perturbation theory of light–matter interaction delivers already the correct scaling factor of the absolute signal strength in terms of an appropriate response function. This is, however, not applicable in our situation because we chose a simplified treatment for reasons of computational speed, rather than dealing with a more realistic but much more complicated full spatial–temporal model [40, 41, 42, 43, 44]. In particular, the response-function treatment provides the field at one particular spatial position only, and one then has to solve the wave equation with that nonlinear field as a source term upon propagation through the full length of the material. This is beyond our scope. Thus, we do not take into account propagation effects and want the final signal to scale simply with the square of the incident field. Then we have to ensure, however, that the pulse energy in the second-harmonic field does not exceed the sum of pulse energies in the initial fields because then energy conservation could not be fulfilled. In practice, we will require only a (small) fraction of the power to be converted in order to prevent saturation effects. We derive in Section LABEL:SISecAbsoluteAmplitudeScalingOfSecondOrderFields of the SI an appropriate scaling factor.
If we allow only a fraction $\eta_{2}$ of the fundamental pulse energy $W_{0}$ to be converted to SHG under optimum conditions (leading to a maximum SHG pulse energy $W_{\text{SHG,max}}$), the final result for the pulse-energy scaling factor is
$S_{\text{SFG}}=\eta_{2}\frac{W_{0}}{W_{\text{SHG,max}}}S_{\text{SFG,unscaled}}.$ (52)
We may set $0<\eta_{2}\leq 1$ at will to guarantee energy conservation in conjunction with an appropriate reduction of the fundamental energies as explained in Section 4.5. We may select, e.g., $\eta_{2}=10\,\mathrm{\%}$ to stay away from the saturation regime. This evaluation has to be done only once for a given laser (unless the laser output parameters are changed). The intermediate multiplication result $\eta_{2}W_{0}/W_{\text{SHG,max}}$ can then simply be stored as a number and used as a multiplier on $S_{\text{SHG,unscaled}}$.
4.4 Nonlinear signal beam curvature
We now discuss the beam radius and radius of curvature of a nonlinear signal beam that arises from the corresponding properties of the incident fields. The incident fields may be mutually different in general, e.g., in pump–probe experiments with different beam radii of pump and probe beams. We thus need to address where the focus of the resulting beam will be located. A detailed derivation is presented in Section LABEL:SISecDerivationOfNonlinearSignalBeamCurvature of the SI, and we only quote the results here.
It turns out that the $n$th-order nonlinear signal beam radius, $w_{\text{s}}(z)$, at the position of a thin sample, $z$, is obtained from the individual beam radii of the $n$ incident beams, $w_{i}(z)$, via
$\frac{1}{w_{\text{s}}^{2}(z)}=\sum_{i=1}^{n}\frac{1}{w_{i}^{2}(z)},$ (53)
in generalization of Eq. (45). The curvature radius of the signal beam at the sample position, $R_{\text{s}}(z)$, is given by
$R_{\text{s}}(z)=\frac{\pi w_{\text{s}}^{2}(z)}{\lambda\zeta_{\text{s}}}$ (54)
with the longitudinal position parameter
$\zeta_{\text{s}}=w_{\text{s}}^{2}(z)\sum_{i=1}^{n}\frac{\alpha_{i}\zeta_{i}}{w_{i}^{2}(z)}$ (55)
with phase-matching coefficients $\alpha_{i}\in\{-1,+1\}$ of the incident beams indicating with which sign they contribute to the calculation of the nonlinear signal phase-matching direction, and
$\zeta_{i}=-\frac{\operatorname{Re}q_{i}}{\operatorname{Im}q_{i}}$ (56)
using the complex beam parameters $q_{i}$ of the incident beams [compare Eq. (LABEL:SIEqComplexRadiusOfCurvature)].
Figure 7: Nonlinear signal beam. (a) Lateral position of the SHG beam waist emerging from two collinear fundamental beams with the same radius while the second beam is laterally shifted. The lateral position of the SHG beam is given relative to that of the first fundamental beam. (b) Beam waist of the SHG beam depending on the beam waist of the second fundamental beam. The waist of the first fundamental beam is set to 1 mm.
For one particular limiting case, consider that the sample or nonlinear crystal is placed near the position of the beam waist of all incident beams, i.e., $\zeta_{i}\ll 1$. This represents the “plane-wave” situation (for which $R_{i}\to\infty$) that can be realized either by placing the sample at the focal point or by employing collimated beams. Then, $\zeta_{\text{s}}\ll 1$, and the signal focus is located at the same longitudinal position as the incident beams are. In other words, the signal beam is created at its waist in the sample.
In Fig. 7(a), we illustrate different focusing conditions of two incoming beams by plotting the lateral SHG beam waist position while varying the lateral displacement between the incoming beams. We consider two collinear beams under perfect phase matching. In case of perfect co-focusing, the waist position of the SHG beam is the same as that of the incident beams. When one of the incident beams is laterally displaced, the SHG beam waist position also shifts, but by half the value of the displacement. For SHG, the Rayleigh range $z_{\text{R}}$ is the same as that of the fundamental [22]. Thus, the half-divergence angle of SHG is smaller by $1/\sqrt{2}$ with respect to the fundamental according to Eq. (LABEL:SIEqHalfDivergenceAngle), with the beam waist smaller by the same factor. If the two beams are co-focused on the sample while the first beam has a fixed radius and the radius of the other beam is increased, the radius of the SHG beam approaches that of the first beam asymptotically [Fig. 7(b)].
Consider for further illustration, as another exemplary special case, pump–probe transient absorption spectroscopy, even though its response-function treatment is beyond the scope of the present work. Nevertheless, this case is instructive for illuminating the generality of the derived beam-parameter results. In that case, the phase-matching direction of the signal is given by
$\mathbf{k}_{\text{s}}=-\mathbf{k}_{\text{pu}}+\mathbf{k}_{\text{pu}}+\mathbf{k}_{\text{pr}}$ (57)
for incident pump ($\mathbf{k}_{\text{pu}}$) and probe ($\mathbf{k}_{\text{pr}}$) wave vectors. The wave vector $\mathbf{k}_{\text{pu}}$ appears twice in this equation, once with a positive sign and once with a negative sign, because there are two (phase-conjugate) interactions with the pump pulse. Thus, the phase-matching coefficients are $\{\alpha_{1},\alpha_{2},\alpha_{3}\}=\{-1,+1,+1\}$ and the signal propagates into the direction of the probe beam (leading to constructive or destructive interference that is then interpreted in measurements as absorbance changes). From Eq. (53), we obtain
$\frac{1}{w_{\text{s}}^{2}(z)}=\frac{2}{w_{\text{pu}}^{2}(z)}+\frac{1}{w_{\text{pr}}^{2}(z)}.$ (58)
It is often customary to choose the pump-beam radius (much) larger than the probe-beam radius. Under this condition, we find from Eq. (58) that $w_{\text{s}}\approx w_{\text{pr}}$ and from Eq. (55) that $\zeta_{\text{s}}\approx\zeta_{\text{pr}}$, and thus the signal beam has the exact same characteristics as the probe beam. This is the reason for choosing a larger pump-beam diameter because then the interference leads to a spatially uniform transient absorption signature across the full probe-beam profile. Otherwise, the signal beam might deviate from the probe beam and one loses contrast.
4.5 Reduction of fundamental pulse energies upon nonlinear signal generation
When new signal fields are generated according to second-order or third-order response and no energy is absorbed in the nonlinear crystal or sample, the total energy summed up over all laser beams should be conserved. In a fully self-consistent solution of Maxwell’s equations, this would arise automatically. Our simplified treatment calls for manual adjustment of fundamental pulse energies. Otherwise, it might happen that the total energy of all output beams after a nonlinear GOE exceeds the total energy of all incident beams.
We arrive at the correct energy, contributed to the new beam by each fundamental, by analyzing the relevant number of photons for each microscopic signal-generation process. Let us consider sum-frequency generation (SFG) as in Eq. (40). Each SFG photon has an energy of $\hbar\omega_{\text{SFG}}$, and the fundamental beams $\tilde{E}_{1}(t)$ and $\tilde{E}_{2}(t)$ carry photons of energy $\hbar\omega_{1}$ and $\hbar\omega_{2}$, respectively. The incident fundamental pulse energies, $W_{1,\text{in}}$ and $W_{2,\text{in}}$, and the generated SFG pulse energy, $W_{\text{SFG}}$, can be obtained from the fields via Eq. (LABEL:SIEqEnergyIsScaleTimesNormSquared) from the SI. Thus, a number of $W_{\text{SFG}}/(\hbar\omega_{\text{SFG}})$ photons have to be generated. This requires the same number of photons contributed from each of the incident beams, reducing their energies accordingly by
$\Delta W_{1}=\frac{\hbar\omega_{1}}{\hbar\omega_{\text{SFG}}}W_{\text{SFG}},\quad\Delta W_{2}=\frac{\hbar\omega_{2}}{\hbar\omega_{\text{SFG}}}W_{\text{SFG}}.$ (59)
Analogously, one obtains the beam reductions for second-harmonic generation by setting $W_{2}=W_{1}$, i.e., the fundamental beam is reduced by twice the value (in addition to the reduction from potential SFG contributions). Furthermore, it is possible that more than two beams interact in any given GOE with second-order response. In that case, we calculate the reductions for all beam-pair combinations (SFG and SHG) and then, in the end, obtain the reduced fundamental output pulse energies by
$W_{1,\text{out}}=W_{1,\text{in}}-\sum_{k}\Delta W_{1,k},\quad W_{2,\text{out}}=W_{2,\text{in}}-\sum_{k}\Delta W_{2,k},$ (60)
where the index $k$ runs over all pairwise combined processes. Note that we reduce the fundamental pulse energies only after having generated all second-order beams with the full incident energy. This ensures that no particular pair of beams is “prioritized” and has a higher pulse energy than any of the others when calculating the SFG and SHG beams successively. Of course the beam-energy reductions occur in addition to those from beam clipping as discussed in Section 3.1, and we do not modify $W$ directly but rather the pulse-energy scaling factor $S$ from Eq. (LABEL:SIEqEnergyIsScaleTimesNormSquared). Working with photon numbers and center frequencies only is an approximation. In principle, one could analyze the complete spectral intensity shape for the correct result [32], but since we require only the adjustment of absolute energies here, such small variations are irrelevant. For the spectral–temporal shapes of the signal and transmitted pulses themselves we take into account the explicit field profiles.
5 Graphical laser beam representation
Simulating linear and nonlinear optical phenomena leads to laser beams propagating through space with parameters as determined in Section 2, Section 3, and Section 4. If one only seeks numerical results, then the calculation is already complete because the obtained parameters provide a full characterization. Often it is desired, however, to represent the laser beams graphically without compromising the real-time simulation. The details of graphical visualization of laser beams will depend on the particular software implementation, which is beyond the scope of the present work. Instead, we discuss some general points that are relevant for any such three-dimensional modeling that is sufficiently fast and realistic.
5.1 Gaussian beam scaling
In this section, we investigate general spatial properties of Gaussian beams. The question will be how the shapes of Gaussian laser beams are related to each other given arbitrary beam parameters. We will find that any Gaussian beam can be derived from the general shape of a “reference beam” by suitable scaling. Thus, it is not necessary to re-calculate, point by point, the spatial envelopes of a beam for any new set of beam parameters, but one can simply generate a Gaussian model once, as will be shown in Section 5.2, and then scale it accordingly.
Let us assume that we start with a suitable representation of a Gaussian “reference beam” that is given by the beam radius,
$\tilde{w}(z)=\tilde{w}_{0}\sqrt{1+\left(\frac{z}{\tilde{z}_{\text{R}}}\right)^{2}},$ (61)
as a function of the distance $z$ from the beam waist with a “Rayleigh length”
$\tilde{z}_{\text{R}}=\frac{\pi\tilde{w}_{0}^{2}}{\lambda M^{2}},$ (62)
for propagation along the $\hat{\mathbf{z}}$ direction, wavelength $\lambda$, and beam-quality parameter $M^{2}\geq 1$ (for Gaussian beams, $M^{2}=1$), with the waist $\tilde{w}_{0}$ located at $z=0$. The tilde symbolizes the “reference” character. We seek a new beam
$w(z)=w_{0}\sqrt{1+\left(\frac{z-z_{0}}{z_{\text{R}}}\right)^{2}}$ (63)
with a new Rayleigh length
$z_{\text{R}}=\frac{\pi w_{0}^{2}}{\lambda M^{2}}$ (64)
whose (potentially different) waist $w_{0}$ is located at position $z=z_{0}$, and we would like to obtain this beam as a suitably transformed version of the reference beam.
It is often convenient to treat Gaussian beam propagation using a complex-valued beam curvature radius $q$, as introduced in Eq. (LABEL:SIEqComplexRadiusOfCurvature) in the SI, that is defined as [45]
$\frac{1}{q(z)}=\frac{1}{R(z)}+i\frac{\lambda M^{2}}{\pi w^{2}(z)}$ (65)
at a position $z$ with relation to the wavefront curvature radius $R$. Beam propagation in free space follows [46, 47]
$q(z)=q_{0}+z$ (66)
for propagation by $z$, and $q_{0}=q(0)$ [see also Eq. (LABEL:SIEqPropagationFreeSpace) in the SI]. Optical elements such as lenses or mirrors will lead to a modification of $q$ [14].
According to Eq. (66), we have
$\tilde{q}(z)=\tilde{q}_{0}+z$ (67)
and
$\tilde{q}_{0}=-i\tilde{z}_{\text{R}}$ (68)
for the propagation law of the reference beam, and its complex curvature at its waist position, respectively. Assume we are given $q=\operatorname{Re}q+i\operatorname{Im}q$, with $\operatorname{Re}q$ and $\operatorname{Im}q$ providing the real and imaginary parts of the complex radius of curvature $q$ of the “new” beam, immediately after transfer through an optical element. Then the new beam waist is located at a position $z_{0}=-\operatorname{Re}q$, translated with respect to the current position at the optical element, and the transformed Rayleigh length is given by $z_{R}=-\operatorname{Im}q$.
Let us define the scaling factor
$b=\frac{z_{\text{R}}}{\tilde{z}_{\text{R}}}=-\frac{\operatorname{Im}q}{\tilde{z}_{\text{R}}}$ (69)
such that
$z_{\text{R}}=b\tilde{z}_{\text{R}}$ (70)
and
$\frac{w_{0}}{\tilde{w}_{0}}=\frac{z_{\text{R}}^{2}}{\tilde{z}_{\text{R}}^{2}}=b^{2}.$ (71)
Then we obtain
$w(z+z_{0})$ $=w_{0}\sqrt{1+\left(\frac{z}{z_{\text{R}}}\right)^{2}}$ (72)
$=b^{2}\tilde{w}_{0}\sqrt{1+\left(\frac{z}{b\tilde{z}_{\text{R}}}\right)^{2}}$ (73)
$=b^{2}\tilde{w}\left(\frac{z}{b}\right).$ (74)
Thus, the new beam $w$ can be obtained from the reference beam $\tilde{w}$ by a scaling factor $b^{2}$ along the transverse coordinates ($x$ and $y$) and by a scaling factor $b$ along the longitudinal propagation coordinate $z$ [noting that the beam is increased in size for $b>1$ along transverse as well as longitudinal coordinates despite the factor $b$ appearing once in the numerator and once in the denominator of Eq. (74)]. The result is then shifted by $-z_{0}$, i.e., by $+\operatorname{Re}q$, along the positive propagation direction.
For representing finite-length beams, one can “clip” the resulting object appropriately at its end positions such that only the relevant section (e.g., between two mirrors) is displayed.
5.2 Three-dimensional beam model
Now that we have seen in Section 5.1 how any Gaussian laser beam can be obtained from one given reference shape, we discuss how such a reference shape can be prepared efficiently. In three-dimensional computer graphics, solid objects are generally modeled as polygon meshes, i.e., a set of points in three-dimensional space (called “vertices”) that are connected by straight lines (“edges”) that in turn form small planes (“faces”). The collection of vertices and lines is sought such that the resulting set of faces approximates the real surface of the given object.
In practice, striving for good performance in real-time rendering applications, one seeks to keep the number of vertices as small as possible to reduce calculation time, yet still capturing the essential shape of the underlying object as accurately as possible. In the present case, we desire to find the minimum number of vertices and their positions that can be used to represent a Gaussian laser beam. The challenge will be that the beam radius of a Gaussian evolves as a hyperbola along the propagation direction as given by Eq. (61) and as a circle around the circumference with radius $w$, while the edges of the mesh are straight line segments. For optimizing the number and distribution of vertices, we introduce a local relative length error, $\varepsilon$, defined as the relative length difference between a curve segment following the real Gaussian shape and a straight edge, connecting the same two neighboring vertices. Then we demand this relative error to be below a certain threshold for all the edges in the mesh, and we seek a vertex distribution that minimizes the number of required vertices for a user-specified error.
We derive in Section LABEL:SISecOptimalVertexDistribution of the SI the minimum number of vertices along the propagation direction for any Gaussian beam, independent of its beam parameters, and find that just five vertices along the propagation direction are sufficient to ascertain a relative maximum error of $\varepsilon=1\,\mathrm{\%}$. This signifies that in many cases, Gaussian beams are well approximated by straight lines at long distances, i.e., the asymptotes of the hyperbolic divergence that is found for $z\gg\tilde{z}_{\text{R}}$ in Eq. (61), and that for the curving sections (between the Rayleigh points) few additional vertices are sufficient. Around the circumference, we find that $13$ vertices are required for $\varepsilon=1\,\mathrm{\%}$, signifying that, compared to the longitudinal direction, more straight edges are required to approximate a circle. Altogether, thus, one needs a minimum of $5\times 13=65$ vertices. Requiring an error of $\varepsilon=0.5\,\mathrm{\%}$ changes these numbers to $7\times 18=126$.
In practice, it may not be practical to use so few vertices along the longitudinal direction because then the resulting mesh might be much larger than the required beam segment between two optical elements. In that case, there would be no mesh vertices at all in the required spatial region. While it is still possible, using suitable rendering agorithms with clipping, to display objects whose vertices are all located outside of the visible region, it may become preferable to define the object with closer-spaced vertices. In particular, the mesh vertices may be required not only for a faithful representation of an (infinitely long) Gaussian beam, but they are also used to terminate the graphical representation when the laser intersects with an object. For that reason, it is helpful to adopt a smaller-step longitudinal mesh spacing in practice by selecting an appropriate maximum step size between any two vertices along the propagation direction, i.e., by selecting an appropriate $\Delta z_{\text{max}}$ in Eq. (LABEL:SIEqOptimalLongitudinalMeshSpacing). The present analysis assures, however, that the relative local error $\varepsilon$ will always be smaller than the given error threshold, and thus the graphical output will approximate the real beam shape well.
5.3 Laser color
When a laser pulse is scattered off a surface, an observer can perceive the spot with a certain color and a certain luminance (“brightness”) that depends on the pulse spectrum, energy, beam radius, and surface properties. For a realistic visualization of such a cross section or of the full laser-beam mesh discussed in Section 5.1 and Section 5.2, we should assign a corresponding computer-graphics color code. For this purpose, we adapt procedures from computer graphics visualization.
In Section LABEL:SISecColorPerception of the SI, we use a “standard observer” model of average human color perception as defined by the “Commission Internationale de l’Éclairage” (CIE). We use the XYZ standard in its most recent implementation [48] and employ the “2-deg XYZ color matching functions transformed from the CIE (2006) 2-deg LMS cone fundamentals” to get the XYZ color values for the given laser spectrum. This color code is subsequently transformed to an RGB representation that can be displayed in computer graphics. In addition, one can use the “alpha” channel A in the RGBA system (typically representing transparency or opacity) to define luminance, as shown in Eq. (LABEL:SIEqLuminance). This luminance is proportional to the pulse-energy scaling factor $S$, meaning a laser beam with higher energy will appear brighter.
Since we account for the full laser spectrum in the color calculation, this automatically leads to the effect that very short pulses (nearing a super-continuum) will appear more “white” compared to narrowband spectra. An adequate representation of the laser color can thus serve a useful function by providing feedback in real-time optics simulations.
6 Simulation results
Figure 8: Frequency-resolved optical gating (FROG) in VR. (a) Screenshot from femtoPro showing a Mach–Zehnder-type interferometer with second-harmonic generation (SHG) using a $\beta$-barium borate (BBO) crystal. (b) Assuming perfect phase matching, the SHG beam in the direction of the superposition of the wave vectors of the two incident beams is detected on a spectrometer. (c) “Experimental” FROG trace obtained in VR. (d) Reconstructed FROG trace. (e) Reconstructed temporal intensity and phase. (f) Reconstructed spectral intensity and phase.
Having introduced our model for linear and nonlinear optics simulations, we now show exemplary simulation results. For that purpose, we consider a setup for a well-known pulse-characterization method, SHG frequency-resolved optical gating (FROG) [49, 26]. The setup as shown in Fig. 8(a) is based on a Mach–Zehnder-type interferometer. Note that, to achieve perfect interferometer alignment, the GOEs shown in Fig. 8(a) were placed programmatically by setting numerical position values rather than by placing them “by hand” in VR. The laser emitter in the top-left corner of Fig. 8(a) generates Gaussian pulses with a center wavelength of 800 nm and $\tau_{\text{p}}=15$ fs duration [intensity full width at half maximum as in Eq. (LABEL:SIEqNonNormalizedTemporalEnvelope)]. After passing an idealized beam splitter without dispersion, the reflected beam with wave vector $\mathbf{k}_{1}$ is delayed using a delay stage which can be controlled “remotely” through a virtual computer interface within the VR environment. The transmitted beam with wave vector $\mathbf{k}_{2}$ is noncollinearly overlapped with the reflected beam in a GOE that has the material parameters of a $\beta$-barium borate (BBO) crystal. Under the assumption of perfect phase matching, the second-order response of the BBO was modeled to be “complete,” i.e., the second harmonics of the two individual beams in $2\mathbf{k}_{1}$ and $2\mathbf{k}_{2}$ direction are generated in addition to the “collaborative” SHG in $\mathbf{k}_{1}+\mathbf{k}_{2}$ direction [Fig. 8(b)]. The latter was detected by a spectrometer. In the scan procedure in VR, the delay stage was incremented from $-100$ fs to $+100$ fs in steps of $0.5$ fs. Rather than directly calculating the resulting FROG trace from an idealized “FROG signal equation,” the data was acquired analogously to reality, i.e., after moving the delay stage to a new step position, the SHG spectrum resulting from the interference of $\mathbf{k}_{1}$ and $\mathbf{k}_{2}$ in the BBO was calculated according to Section 4.3 before moving to the next step position.
The entire scan took $22\,\mathrm{s}$, and the consistently high and stable frame rate allowed real-time monitoring of changes in the fringe pattern detected by the spectrometer at each scanning step. The thus-obtained “experimental” FROG trace shown in Fig. 8(c) was evaluated using a commercial FROG analysis program [50], yielding the reconstruction results shown in Fig. 8(d–f).
Using a grid size of 64, the minimum FROG error was $2.5\times 10^{-4}$. Since simulated data have been used, the reconstruction quality is excellent [compare Fig. 8(c) and Fig. 8(d)]. For a more immersive experience, artificial noise could be added to the acquired data. The slight up-chirp observed in the reconstructed electric field is a result of phase accumulation by propagation through air over the distance from the laser emitter to the BBO, leading to a reconstructed pulse duration of $17.4$ fs. Note that we use the approximation of a thin BBO, i.e., the SHG signal is calculated at the front surface of the BBO crystal such that further dispersion introduced by the finite thickness of the BBO is not taken into account.
7 Conclusion
In the present work, we have described the development of a model for real-time simulation of linear and nonlinear optical phenomena and spectroscopy. The term “real time” means that we endeavor to find such a level of approximation that allows calculation times of complete experimental configurations on the order of $10\,\mathrm{ms}$ on consumer-grade hardware. We chose to implement a mixture of geometrical optics and Gaussian wave optics because the former offers computational speed and simplicity, while the latter provides the correct evolution of beam radius and beam curvature during propagation. A finite beam radius is in turn important to describe effects of beam overlap in interference phenomena and nonlinear signal generation. A summary of features and limitations is given in tabular form in Section LABEL:SISecFeaturesAndLimitations of the SI.
In particular, we provided expressions that can be used to computationally process frequency- and time-sampled electric field evolutions of femtosecond laser pulses. We removed fast phase oscillations due to the carrier frequency (from the time-domain field) and due to propagation (from the frequency-domain field) to facilitate numerical stability. In addition, a suitable normalization condition separates the pulse energy and the amplitude changes that occur due to beam-radius evolution.
We derived equations that can be used to calculate geometric beam parameters after transmission of laser beams through finite circular apertures, ignoring astigmatism and diffraction. The obtained approximations are straightforward to apply and ensure that geometric laser-beam cross sections are always contained within the GOE apertures from which they emerge. The energy throughput was determined from the numerical evaluation of a Gaussian transmission integral that was implemented as a precalculated lookup table.
Considering multi-beam linear interference, we obtained a contrast visibility factor that takes into account the individual parameters of each incident beam.
Thus, we treated the case that an arbitrary number of laser beams – which arrive from different directions, are laterally displaced, and have different individual amplitudes and beam radii – are superimposed on a detector plane. While not resolving the detailed spatial interference pattern, we obtained integrated results as measured by a spectrometer or power meter.
Laser pulses are modified by passing through GOEs. In particular, we considered the response-function formalism to describe linear non-resonant dispersion, resonant absorption (using a Franck–Condon model), and second-order non-resonant response. The latter allowed inclusion of sum-frequency and second-harmonic generation. While such processes had been treated extensively in the scientific literature, we derived simplified approximative results that nevertheless take into account the Gaussian beam overlap between generating pulses and the Gaussian beam curvatures (i.e., focusing properties), and how both properties translate to the nonlinearly generated signal beam.
We derived general spatial properties of Gaussian laser beams that are relevant for graphical representation. In particular, we found that the spatial envelope of any Gaussian beam can be visualized, at a $1\,\mathrm{\%}$ error level, in three-dimensional space using a polygonal mesh with just five vertices connected by straight edges along the propagation direction to describe the hyperbolic beam evolution and 13 vertices along the circumference to approximate the circular symmetry. Such a rendered object can then be scaled to represent Gaussian beams with any other spatial parameters, and we derived the scaling factors. For color visualization, we also obtained the relation between the physical laser spectrum and the computer-graphics RGB color system.
The model described in the current work forms the core of an interactive and immersive virtual-reality (VR) simulation of an ultrafast laser laboratory that we have recently developed [14, 15]. The results of the present work are also applicable, though, for other optics or spectroscopy simulations outside of femtoPro and beyond a VR context.
[…]
Funding This work was supported by the Fonds der Chemischen Industrie (FCI). We further acknowledge funding by the Julius-Maximilians-Universität Würzburg Project WueDive via Stiftung Innovation in der Hochschullehre.
[…]
Acknowledgment We thank Pavel Malý for deriving equations of Gaussian field overlap and nonlinear signal-beam curvature.
[…]
Disclosures “femtoPro” is a registered trademark of the University of Würzburg. The software is available via the femtoPro website [15].
[…]
Data Availability Statement The data that support the findings of this study are openly available at the following DOI: [link will be inserted upon publication].
[…]
Supplemental document See the Supplementary Information for supporting content.

### v2

@@
 While our motivation for the present work thus stems from implementing this VR simulator, the derived equations are much more general and are applicable in other optics simulation scenarios, outside of a VR context, particularly if fast computation is desired.
-Thus, the purpose of the present paper is two-fold: 1) to explain the assumptions underlying the femtoPro software such that users may understand its features and limitations, and 2) to introduce and discuss physical models that can be employed for fast simulations on any platform and for various purposes.
-The requirement for a lag-free user experience sets a top priority on the speed of the underlying simulations.
+Thus, the purpose of the present paper is two-fold: 1) to explain the assumptions underlying the femtoPro software such that users may understand its features and limitations, and 2) to introduce and discuss physical models that can be employed for fast optics simulations on any platform and for various purposes.
+The requirement for a lag-free user experience sets a top priority on the speed of the simulations.
 This is different from the established simulation tools and protocols listed above.
 Still, the accuracy needs to be sufficient to provide a realistic simulation of the spatial and temporal propagation of ultrafast laser pulses and their interaction with matter.
-Some requirements are therefore: We desire to take into account Gaussian beam propagation because finite-focus-size effects are relevant to simulate beam overlap alignment procedures; laser beams may be clipped if they hit the edge of an optical element or an iris; femtosecond pulses are modified by dispersion upon propagation through materials and even more so upon resonant interactions; and second-order processes shall be included to simulate frequency conversion, all in real time as defined above.
+Some requirements are therefore: We desire to take into account Gaussian beam propagation because finite-focus-size effects are relevant to simulate beam overlap alignment procedures; laser beams may be clipped if they hit the edge of an optical element or an iris; femtosecond pulses are modified by dispersion upon propagation through materials and upon resonant interactions; and second-order processes shall be included to simulate frequency conversion, all in real time as defined above.
 The underlying simulation concepts are in principle known from the vast existing scientific literature, in particular from textbooks on general optics [18, 19, 20, 21, 22, 23] and ultrafast spectroscopy [24, 25, 26, 27, 28, 29, 30, 31, 32].
 However, the real-time simulation requirement necessitates a careful balance between speed and accuracy.
-In the present work, we derive simplified expressions that facilitate real-time optics simulations.
-We start by defining our representation of the electric fields of laser pulses (Section 2).
-Then we discuss the spatial modulations of laser beams resulting from geometric effects of optical elements and multi-beam interference (Section 3), followed by spatial and spectral–temporal modulations arising from linear and nonlinear light–matter interaction (Section 4) and the graphical representation of laser beams (Section 5).
+In another manuscript, we describe the concept, implementation, and evaluation of a dynamic graph-based algorithm that is used to find a self-consistent solution of multiple propagating and interfering laser beams [33].
+The dynamic graph model maps the formal network between beam segments of laser pulses propagating freely in space as “edges,” and optical elements (such as mirrors or lenses) that modify the laser pulse properties as “nodes,” and then the algorithm is set up to minimize the number of evaluation steps.
+In the present work, we focus on the physical model underlying the individual calculation steps of that algorithm and derive simplified mathematical expressions that facilitate real-time optics simulations.
+We start by defining our representation of the electric fields of laser pulses including the spatial modulations of laser beams resulting from geometric effects of optical elements (Section 2).
+Then we discuss multi-beam interference (Section 3), followed by spatial and spectral–temporal modulations arising from linear and nonlinear light–matter interaction (Section 4) and the graphical representation of laser beams (Section 5).
 Exemplary simulation results are shown in Section 6 before we conclude in Section 7.
-2 Laser pulses
-The representation of electric fields of laser pulses is treated in many comprehensive books on general optics [18, 19, 20, 21, 22, 23] and ultrafast time-resolved spectroscopy [24, 25, 26, 27, 28, 29, 30, 31, 32].
-Here we seek a representation that is suitable for efficient calculations in computational models.
-Since we need to describe both spatial and spectral–temporal properties, we have to separate the various aspects in a convenient manner.
-We define and derive all relevant expressions in great detail in Section LABEL:SISecElectricField of Supplementary Information (SI) and only quote the main results here.
+2 Laser pulse propagation
+We define and derive all relevant expressions for the description of laser pulses in great detail in Section LABEL:SISecElectricField of Supplementary Information (SI) and only quote the main results here.
+Indeed, we highly recommend reading the entire Section LABEL:SISecLaserPulsePropagationAndModificationByOpticalElements of the SI before continuing with the main article here because all the relevant quantities and essential transformations are set up in the SI.
 As a result of the derivations, we find that the electric field of a laser pulse at position $\mathbf{r}$ as a function of spatial coordinates $x$, $y$ and $z$ and of time $t$ is given by
@@
 (LABEL:SIEqEPropXYZT) in the SI, where the “$+$” superscript in $E_{\rm{prop}}^{+}$ indicates that this is a complex-valued representation resulting from the positive-frequency part of the full field only and the subscript “prop” indicates that the pulse has propagated through space along the $\hat{\mathbf{z}}$ direction requiring the propagation time $T$.
-The beam radius at each $z$ position is given by $w(z)$ (with details on the evolution of $w(z)$ as a function of $z$ provided in Section 5) and lateral translations of the beam along $\hat{\mathbf{x}}$ and $\hat{\mathbf{y}}$ by $x_{0}$ and $y_{0}$, respectively.
+The beam radius at each $z$ position is given by $w(z)$ [Eq.
+(LABEL:SIEqBeamRadiusEvolution)] and lateral translations of the beam along $\hat{\mathbf{x}}$ and $\hat{\mathbf{y}}$ by $x_{0}$ and $y_{0}$, respectively.
 Further, $\mathbf{k}$ denotes the wave vector, $\omega_{0}$ the center frequency, $\delta t$ the sampling step size in time domain, $S$ a scaling factor proportional to pulse energy, and $\tilde{E}_{t}(j)$ the numerical array representation of the complex-valued temporal envelope (including amplitude and phase terms) for individual sampling points that are indexed by $j$.
@@
 with analogous definitions, and where $\delta\omega$ describes the sampling step size in frequency domain.
-3 Geometric effects
-3.1 Finite aperture
+We now discuss how to describe the propagation of laser beams in the presence of finite apertures.
 Apertures lead to the clipping of laser beams.
@@
 Thus, in the present work we suggest an intermediate regime where we treat some phenomena in the limit of geometrical optics and some phenomena using wave optics.
-Concerning apertures, we select a treatment by geometrical optics.
+Concerning apertures, we select a treatment by geometrical optics as discussed in detail in Section LABEL:SISecFiniteAperture of the SI.
 This has the advantage that we do not have to follow various diffracted beamlets that might propagate in various different directions, in particular after interaction with subsequent optical elements.
 Instead, we can continue to describe the transmitted beam as a single entity with one limited set of parameters.
-While this is a limitation that removes some phenomena from being treated with the model, it strongly reduces complexity and facilitates real-time simulations.
-Despite this approximation, we seek a representation that is as faithful to reality as possible under this approximation.
+The approximation of geometrical optics entails that our model cannot be used to simulate diffraction gratings.
+On the other hand, this strongly reduces complexity and facilitates real-time simulations.
+We seek a representation that is as faithful to reality as possible under the given model assumptions.
 Thus, a laser beam emerging from any open aperture should not have a cross section that extends beyond the hard limits of the aperture.
@@
 The laser beam then would be centered onto at least two subsequent irises, and after the alignment procedure is completed, the irises would be opened to let the full beam pass.
-In that case, then, the final beam propagation would again not be limited by diffraction for the “real” experiment, justifying the limit of geometrical optics, while the clipping effects will be captured at least qualitatively or semi-quantitatively (to be discussed below) during the simulation of alignment procedures.
+In that case, then, the final beam propagation would again not be limited by diffraction for the “real” experiment, justifying the limit of geometrical optics, while the clipping effects will be captured at least qualitatively or semi-quantitatively during the simulation of alignment procedures.
 As a second possibility, finite-aperture effects become relevant when a laser beam is larger than the optical element it hits, or when it hits the element at its edge.
@@
 Thus, again, it seems justified to employ a simplified geometrical optics treatment.
-The approximation of geometrical optics entails that our model cannot be used to simulate diffraction gratings.
-This is a current limitation (see Section LABEL:SISecFeaturesAndLimitations in the SI for an overview of features and limitations).
-In order to ensure that the Gaussian beam that is transmitted through an aperture does not have a cross section that extends beyond the aperture limits, one has to evaluate the geometrical common cross section of the laser beam and the aperture.
-At normal incidence, one requires the common area of two circles formed by the aperture and the laser cross section, in which one then has to fit a third circle that does not extend beyond the common limits.
-For nonlinear signal generation from at least two separate beams, however, one would already have to take into account at least three circles, i.e., one for the aperture and two for the incident beams.
-The problem of the common intersection of three circles is already quite involved with an analytical solution reported only in 2006 [34].
-Anyway, the laser beam will, in general, not hit the aperture at normal incidence.
-In that case, one would have to consider the common overlap of three ellipses that is even more complicated than that of three circles because of the many types of overlap cases that have to be distinguished [35, 36].
-Since we ignore astigmatism, however, it is enough to consider a suitably adapted scenario with circular symmetry.
-We have adopted the “mixed” solution that evaluates overlaps between several beams, such as in interference evaluation or nonlinear signal generation, on the basis of overlapping Gaussian functions; as a second component of our approach, we then treat the transmission through a hard aperture in a purely geometrical fashion to ensure that “no” laser beam profile extends beyond the open aperture.
-The term “no” is set in quotation marks here because the visual appearance of the laser beam, when rendered as a three-dimensional object up to its beam radius, will indeed not be larger than the transmission aperture.
-On the other hand, since we treat the beam as a Gaussian, this still allows its radial distribution to extend further outwards than the aperture opening.
-This is a consequence of the hybrid approach, adopted in this work, between geometrical optics and wave optics when diffraction is ignored.
-Thus, we take the beam radius as the geometric limit.
-When we then choose to visualize laser beams via plotting their beam radius, an intuitively consistent behavior emerges if the beam waist is taken to represent the “edge.”
 For treating the transmission through optics, we have to consider finite incidence angles.
 For example, mirrors are routinely employed at $\approx 45^{\circ}$ incidence angle to deflect beams, and anyway we cannot avoid that the beam hits a general optical element (GOE) at an arbitrary angle, so the model needs to deal with all situations.
-Thus we may not assume the “paraxial” limit, and we employ vector calculus for an analysis independent of any particular chosen coordinate system.
-The relevant geometry is illustrated in Fig.
-1.
-The incident laser beam (from the left, potentially arriving at an additional angle elevated out of the drawing plane) is given by a point of origin in world coordinates, $\mathbf{r}_{\text{in}}$, in our evaluation logic usually its starting point at the previous GOE; a unit directional vector $\hat{\mathbf{k}}_{\text{in}}=\mathbf{k}_{\text{in}}/\left\lvert\mathbf{k}_{\text{in}}\right\rvert$, where $\mathbf{k}_{\text{in}}$ is the incident wave vector; and a beam radius $w_{\text{in}}$ at the point of intersection with the current GOE.
-The current GOE is defined by its aperture center in world coordinates, $\mathbf{o}$, a normal vector describing the orientation of the “active” aperture plane, $\hat{\mathbf{n}}$, and an aperture radius $a$.
-Figure 1: Aperture geometry.
-The input beam (red) propagates along a central “ray” (solid) with lateral width determined by its beam radius (dashed–dotted) until it intersects at an incidence angle $\alpha$ with the central active plane of a general optical element (GOE) that is assumed, without loss of generality, to be normal to the “optical axis” (dashed black).
-The output beam (green) is likewise defined by a central ray (solid) and radius (dashed–dotted).
-All relevant scalar and vectorial quantities are marked in blue as defined in the text.
-We first determine the propagation vector, $\mathbf{L}$, that extends from the origin of the laser beam at the previous GOE to the intersection with the active plane of the current GOE [37],
-$\mathbf{L}=\frac{(\mathbf{o}-\mathbf{r}_{\text{in}})\cdot\hat{\mathbf{n}}}{\hat{\mathbf{k}}_{\text{in}}\cdot\hat{\mathbf{n}}}\,\hat{\mathbf{k}}_{\text{in}}\quad\text{if }\hat{\mathbf{k}}_{\text{in}}\cdot\hat{\mathbf{n}}\neq 0.$ (3)
-If $\hat{\mathbf{k}}_{\text{in}}\cdot\hat{\mathbf{n}}=0$, the laser beam propagates parallel to the GOE’s active plane, and we have to distinguish further cases.
-If, in that case, the numerator in Eq.
-(3) is not equal to zero, the laser beam is parallel to but outside of the surface, so that there is no intersection.
-If the numerator and the denominator in Eq.
-(3) are equal to zero, then the laser beam propagates within the surface, i.e., the GOE is hit exactly from the side.
-We then define the propagation vector such that it propagates the laser beam from its origin to the location closest to the center of the GOE,
-$\mathbf{L}=\mathbf{o}-\mathbf{r}_{\text{in}}-\text{proj}_{\hat{\mathbf{k}}_{\text{in}}}(\mathbf{o}-\mathbf{r}_{\text{in}}),$ (4)
-with the vector projection
-$\text{proj}_{\mathbf{b}}\mathbf{a}=\left(\mathbf{a}\cdot\frac{\mathbf{b}}{\left\lvert\mathbf{b}\right\rvert}\right)\frac{\mathbf{b}}{\left\lvert\mathbf{b}\right\rvert}$ (5)
-of vector $\mathbf{a}$ onto vector $\mathbf{b}$.
-The propagation length $L=\left\lvert\mathbf{L}\right\rvert$ enters Eq.
-(LABEL:SIEqPropagationTime) in the SI for the calculation of the propagation time $T$.
-From the previous origin of the laser and the propagation vector we obtain the “hit” point of intersection with the current GOE, in world coordinates, as
-$\mathbf{h}=\mathbf{r}_{\text{in}}+\mathbf{L},$ (6)
-and as
-$\mathbf{d}=\mathbf{h}-\mathbf{o}$ (7)
-in a GOE-local coordinate system with the origin at the center of the GOE, where we further define the scalar intersection distance $d=\left\lvert\mathbf{d}\right\rvert$ from the GOE center.
-The vectorial quantity $\mathbf{d}$ enters the calculation of directional change by a GOE with focal length not equal to zero, as will be derived below.
-In the present context, we project the incident beam radius, $w_{\text{in}}$, along direction $\mathbf{d}$ to obtain the projected beam radius, $b_{\text{in}}$.
-For this, we observe in Fig.
-1 that $\cos\alpha=w_{\text{in}}/b_{\text{in}}$ but also that $\cos\alpha=\hat{\mathbf{n}}\cdot\hat{\mathbf{k}}_{\text{in}}$.
-Eliminating $\alpha$ provides
-$b_{\text{in}}=\frac{w_{\text{in}}}{\left\lvert\hat{\mathbf{n}}\cdot\hat{\mathbf{k}}_{\text{in}}\right\rvert},$ (8)
-where we introduced the absolute magnitude operator such that it works also if the surface normal $\hat{\mathbf{n}}$ is inverted.
-In the case of $\hat{\mathbf{n}}\cdot\hat{\mathbf{k}}_{\text{in}}=0$, as noted above, the beam travels within the plane, and the projected radius is infinite as it should be.
-This leads to the further treatment as case 2 in the list below.
-Now we have obtained the three quantities, aperture radius $a$, intersection distance $d$, and projected beam radius $b_{\text{in}}$, that are required for the following case distinction based on their quantitative relation:
-  1. 1.
-    $a\geq d+b_{\text{in}}$: The laser beam is completely within the aperture;
-  2. 2.
-    $b_{\text{in}}>d+a$: The laser beam overfills completely the aperture;
-  3. 3.
-    $d>b_{\text{in}}+a$: The laser beam is completely outside the aperture;
-  4. 4.
-    Else: The laser beam is partially within the aperture.
-Let us evaluate, for each case, the projected radius of the new beam after the aperture, $b_{\text{out}}$, along direction $\mathbf{d}$; the origin of the new beam after the aperture in world coordinates, $\mathbf{r}_{\text{out}}$; and the direction of the new beam after the aperture, $\hat{\mathbf{k}}_{\text{out}}$.
-Finally, we obtain the new beam radius perpendicular to the new direction of propagation, $w_{\text{out}}$, that is given in all cases, analogous to Eq.
-(8), by
-$w_{\text{out}}=b_{\text{out}}\left\lvert\hat{\mathbf{n}}\cdot\hat{\mathbf{k}}_{\text{out}}\right\rvert.$ (9)
-Note that in such a treatment, we consider the beam radius as projected along direction $\mathbf{d}$ only, and we ignore the shape and extension in the orthogonal direction.
-This is an approximation that works well for meridional rays but less so for skew rays.
-Nevertheless, we follow this approach because of our choice to ignore astigmatism for computational efficiency [14], and thus one may define just one particular projection coordinate (not two) to represent a beam radius and shape.
-Using the meridional-ray approximation leads to the desired effect that no transmitted beam is larger than any circular open aperture.
-Case 1 ($a\geq d+b_{\text{in}}$) in the list above corresponds to the “normal” alignment goal of having an unclipped beam with an open iris aperture (or a fully reflected beam off a mirror).
-Then the starting point of the new beam is equal to the intersection point of the incident beam,
-$\mathbf{r}_{\text{out}}=\mathbf{h},$ (10)
-the new beam direction is given by [38]
-$\hat{\mathbf{k}}_{\text{out}}=\frac{\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{d}}{f}}{\left\lvert\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{d}}{f}\right\rvert},$ (11)
-and the projected radius is simply
-$b_{\text{out}}=b_{\text{in}}.$ (12)
-In case 2 ($b_{\text{in}}>d+a$), the clipped beam passes, symmetrically distributed, exactly through the center of the GOE, independently of the local intersection coordinate $\mathbf{d}$ that was calculated for the center of the unclipped beam.
-Note that in this geometric treatment, as motivated at the beginning of this section, we do not take into account the Gaussian variation of the intensity within the lateral cross section, and thus we arrive at the statement of the symmetric distribution around the GOE center $\mathbf{o}$.
-Thus, this case corresponds to a “center ray” in geometrical optics that is unaffected by any focal length, and hence we obtain
-$\mathbf{r}_{\text{out}}=\mathbf{o}$ (13)
-for the origin,
-$\hat{\mathbf{k}}_{\text{out}}=\hat{\mathbf{k}}_{\text{in}}$ (14)
-for the direction, and
-$b_{\text{out}}=a$ (15)
-for the projected radius of the new beam.
-Case 3 ($d>b_{\text{in}}+a$) is simple to treat because then no beam is transferred.
-Finally, in case 4 (“else”), the beam is clipped partially. From Fig. 1, one finds $2b_{\text{out}}=a-d+b_{\text{in}}$ such that the projected radius is
-$b_{\text{out}}=\frac{1}{2}(a-d+b_{\text{in}}).$ (16)
-If the beam shall indeed arise from the common overlap between original projected radius and aperture radius, we have to laterally “move” the new beam’s origin such that it is in the center of the new beam.
-From Fig.
-1, one obtains $p=\left\lvert\mathbf{p}\right\rvert=d-b_{\text{in}}+b_{\text{out}}=\frac{1}{2}(a+d-b_{\text{in}})$ and, as a GOE-local vector,
-$\mathbf{p}=p\frac{\mathbf{d}}{d}=\frac{1}{2}(d+a-b_{\text{in}})\frac{\mathbf{d}}{d}$ (17)
-such that
-$\mathbf{r}_{\text{out}}=\mathbf{o}+\mathbf{p}$ (18)
-and
-$\hat{\mathbf{k}}_{\text{out}}=\frac{\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{p}}{f}}{\left\lvert\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{p}}{f}\right\rvert}.$ (19)
-Note that now the effective “distance” entering in Eq.
-(19) is $\mathbf{p}$ rather than $\mathbf{d}$ as in Eq.
-(11).
-We illustrate the case of partial beam clipping in Fig.
-2, where we vary the lateral shift of a beam incident on a GOE with a finite aperture while displaying the waist and the angle of an output beam.
-The radii of the input beam and the aperture are set to 3 mm.
-To invoke a change of the angle of the output beam, we set the focal length of the GOE to $f=37.5$ cm.
-We display two subcases.
-In case of a small angle of incidence (5.71), the output beam waist decreases linearly when shifting the input beam away from the center of the aperture [Fig.
-2(a)].
-The output beam angle also changes linearly.
-In case of a steeper angle of incidence (45) as shown in Fig.
-2(b), the radius and angle of the outgoing beam remain constant in a region around the center of the aperture.
-When the incident beam is further laterally shifted, clipping again results in a change of radius and angle of the outgoing beam.
-Figure 2: Output beam from a finite aperture.
-The lateral shift of an input beam relative to the optical axis is varied under a given incidence angle.
-The beam waist of the input beam and the aperture radius are both chosen to be 3 mm while the focal length of the GOE is set to 37.5 cm.
-The beam waist and the angle of the output beam is shown for an incidence angle of (a) 5.71 and (b) 45.
-The incidence angles are indicated as horizontal dashed lines.
-Finally, we need to consider the reduction in pulse energy due to beam clipping caused by the finite aperture.
-Above, we made the approximation of uniform spatial profiles for generating the geometric parameters of the new beam.
-Let us first consider what would happen at the same level of approximation for the transmitted pulse energy (which, as we will see, is not what we implemented).
-In this simplest approximation of a “flat-top” beam, we could relate the output energy, $W_{\text{out}}$, to the incident pulse energy, $W_{\text{in}}$, by considering the ratio of the projected cross-sectional areas at the aperture,
-$W_{\text{out}}=\frac{\pi b_{\text{out}}^{2}}{\pi b_{\text{in}}^{2}}W_{\text{in}}.$ (20)
-Note that one should not use the orthogonal beam radii $w_{\text{in}}$ and $w_{\text{out}}$ in this context because they are different already if the output beam direction $\hat{\mathbf{k}}_{\text{out}}$ is different from the incidence beam’s direction $\hat{\mathbf{k}}_{\text{in}}$ due to an off-axis focusing element with focal length $f\neq 0$, i.e., the pulse energy does not change but only the intensity changes due to the changed area.
-While Eq.
-(20) describes the situation for a uniform spatial intensity profile, this does not correspond to the properties of a Gaussian beam in which the intensity is higher in the center than at the edge and which is what we introduced in Eq.
-(1) and Eq.
-(2).
-The discrepancy would be noticeable and is relevant if the simulation model is used, for example, for didactic training in the context of beam alignment onto an aperture.
-In those cases, the goal is to align the position of a laser beam with the center of a given aperture (or vice versa, center an aperture on a given laser beam) by monitoring the energy changes detected with a power meter after the aperture when moving the beam (or aperture) laterally.
-Assume now that the laser overfills the aperture (case 2 in the list above).
-In that case, the transmitted cross-sectional area would not change, and neither would the transmitted pulse energy according to simplistic Eq.
-(20).
-Thus, we seek a more accurate calculation of the transmitted energy.
-More accurately, we might alternatively consider, instead of Eq.
-(20) from above, equation (14) from reference [39] that derives, for normal incidence at which $b_{\text{in}}=w$, the following relation:
-$W_{\text{out}}=W_{\text{in}}\exp\left(-\frac{2d^{2}}{w^{2}}\right)\times\sum_{k=0}^{\infty}\left\{\frac{2^{k}d^{2k}}{w^{2k}k!}\left[1-\exp\left(-\frac{2a^{2}}{w^{2}}\right)\sum_{i=0}^{k}\frac{2^{i}a^{2i}}{w^{2i}i!}\right]\right\}.$ (21)
-It is discussed in reference [39] that for central beams, i.e., $d=0$, one obtains the more well-known relation
-$W_{\text{out}}(d=0)=W_{\text{in}}\left[1-\exp\left(-\frac{2a^{2}}{w^{2}}\right)\right],$ (22)
-but this, on its own, is not a useful limit for us either because we want to deal with the particular situation of laterally displaced beams.
-Thus we evaluate the transmitted energy numerically using precalculated lookup tables as derived in Section LABEL:SISecGaussianTransmissionThroughCircularAperture in the SI.
-This allows us to assign transmission values fast for any given combination of beam radius $w$, aperture radius $a$, and displacement $d$.
-Note that we do not change the pulse energy directly, but rather the pulse-energy scaling factor $S$ from Eq.
-(LABEL:SIEqEnergyIsScaleTimesNormSquared).
-The scaling factor as a function of the aperture radius is displayed in Fig.
-3(a), where the behavior of Gaussian beams is evident by an $S$ less than one in the case of $w=a$.
-In Fig.
-3(b), we laterally shift the input beam relative to an aperture with size $a=3w$.
-Note that we capture the beams geometrically up to their beam radius, i.e., we treat Gaussian beams, which theoretically go to infinity, as hard-cut structures.
-This means that if the beam is shifted away from the edge of the aperture by a distance larger than $w$, it is treated as no longer hitting the aperture, leading to zero energy which in turn results in the small “step” visible in Fig.
-3(b) at a lateral shift of $4w$.
-This provides a quantitative estimate for the error at the chosen level of approximation of our model.
-Figure 3: Pulse-energy scaling factor S of a Gaussian beam incident on an aperture.
-(a) Scaling factor as a function of the aperture size.
-(b) Scaling factor as a function of the lateral shift relative to an aperture with a size of $3w$.
-3.2 N-beam interference
+Thus we may not assume the paraxial limit, and we employ vector calculus for an analysis independent of any particular chosen coordinate system.
+In the case of an unclipped beam transmitted through a GOE with focal length $f$, a beam with normalized incident wave vector $\hat{\mathbf{k}}_{\text{in}}$ that hits the GOE at a displacement $d$ with respect to the GOE center will exit the element in a new direction given by the normalized outgoing wave vector $\hat{\mathbf{k}}_{\text{in}}$, where these quantities are related by [35]
+$\hat{\mathbf{k}}_{\text{out}}=\frac{\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{d}}{f}}{\left\lvert\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{d}}{f}\right\rvert}$ (3)
+as derived in Eq. (LABEL:SIEqDirectionChangeGOELaserCompletelyWithinAperture). The beam radius projected onto the plane of the GOE remains the same because there is no clipping.
+In the case of partial beam clipping, the quantity $\mathbf{d}$ in Eq.
+(3) needs to be replaced by an effective quantity $\mathbf{p}$ as in Eq.
+(LABEL:SIEqDefinitionP) that points from the GOE center to the center of the emerging clipped beam, leading to a modified direction of the outgoing beam according to Eq.
+(LABEL:SIEqDirectionChangeGOELaserPartiallyWithinAperture).
+In the case that the beam is fully clipped, the center of the transmitted beam will be at the center of the GOE even if the incident beam was laterally displaced.
+This is because we consider only circularly symmetric apertures such that emerging beams that are fully defined by the aperture will always be centered.
+In that case, the outgoing beam direction is the same as the ingoing beam direction according to Eq.
+(LABEL:SIEqKOutCompleteClipping), which can be rationalized because the direction of a “center beam” through a lens is unaffected by any focal length.
+Finally, we need to consider the power transmission through a circular aperture.
+In Section LABEL:SISecFiniteAperture of the SI, we assume without loss of generality that the beam is displaced along a GOE-local $\hat{\mathbf{x}}$ coordinate with respect to the center of the partially closed GOE aperture.
+Then we have to solve Eq.
+(LABEL:SIEqPowerTransmission),
+$P_{\text{out}}=I_{0}\underset{(x^{2}+y^{2}\leq a^{2})}{\int\mathrm{d}x\int\mathrm{d}y}\,\exp\left[-2\frac{(x-d)^{2}}{w^{2}}-2\frac{y^{2}}{w^{2}}\right],$ (4)
+where the output power $P_{\text{out}}$ is obtained from the on-axis incident beam intensity $I_{0}$, the incident beam radius $w$, the displacement $d$, and the aperture radius $a$ through integration of the spatial Gaussian beam intensity distribution over the circular opening, which is performed numerically with pre-calculated loop-up tables to speed up the computation.
+In the case of a straight edge instead of a circular aperture, the power transmission is obtained in closed form in Eq.
+(LABEL:SIEqTransmissionStraightEdge) via the complementary error function in Eq.
+(LABEL:SIEqComplementaryErrorFunction).
+3 N-beam interference
 We now analyze the interference of $N$ laser pulses, $E_{k,\text{prop}}^{+}(\mathbf{r},\omega)$, $k=\{1,\ldots,N\}$, each given by Eq.
@@
 We proceed in the coordinate system in which the $\hat{\mathbf{z}}$ axis is parallel to the average of the incident wave vectors.
-As will be derived below, interference is visible only for small phase mismatch, and thus the approximation is valid that the beam positions, $(x_{k},y_{k})$ and $(x_{l},y_{l})$, and radii, $w_{k}$ and $w_{l}$, need not be transformed because the average direction is almost identical to either of the incident directions.
+As will be derived in the SI, interference is visible only for small phase mismatch, and thus the approximation is valid that the beam positions, $(x_{k},y_{k})$ and $(x_{l},y_{l})$, and radii, $w_{k}$ and $w_{l}$, need not be transformed because the average direction is almost identical to either of the incident directions.
 The actual incidence angle onto the detector plane is not decisive in first order because the underlying spatial interference pattern is retained, and thus integration over the tilted pattern provides a comparable total power as integration over the normal-incidence pattern.
@@
 This integral is solved in Section LABEL:SISecInterferenceVisibilityFactor of the SI, leading to the already reported result [14]
-$\begin{split}\eta_{k,l}=&\frac{2w_{k}w_{l}}{w_{k}^{2}+w_{l}^{2}}\exp\left[-\frac{(x_{k}-x_{l})^{2}+(y_{k}-y_{l})^{2}}{w_{k}^{2}+w_{l}^{2}}\right]\\ &\times e^{i(\Delta k_{x}x_{\text{p}}+\Delta k_{y}y_{\text{p}})}\exp\left[-\frac{\Delta k_{x}^{2}+\Delta k_{y}^{2}}{2\left(\frac{1}{w_{k}^{2}}+\frac{1}{w_{l}^{2}}\right)}\right].\end{split}$ (28)
+$\eta_{k,l}=\frac{2w_{k}w_{l}}{w_{k}^{2}+w_{l}^{2}}\exp\left[-\frac{(x_{k}-x_{l})^{2}+(y_{k}-y_{l})^{2}}{w_{k}^{2}+w_{l}^{2}}\right]e^{i(\Delta k_{x}x_{\text{p}}+\Delta k_{y}y_{\text{p}})}\exp\left[-\frac{\Delta k_{x}^{2}+\Delta k_{y}^{2}}{2\left(\frac{1}{w_{k}^{2}}+\frac{1}{w_{l}^{2}}\right)}\right].$ (10)
 Note that interference visibility in the literature is often treated in terms of the coherence functions of the light fields.
@@
 For the spectrally integrated result, one then obtains the usual limit of optical coherence and interference visibility that disappears when fields are separated longitudinally by a distance larger than their coherence length (or, for bandwidth-limited ultrashort pulses, by delay times larger than their pulse duration).
-In Fig.
-4(a), we illustrate the dependence of the interference visibility $\eta_{1,2}$ between two beams on various parameters.
-In the first example [Fig.
-4(a)], one beam remains constant and the second beam is shifted laterally with otherwise identical parameters.
-We show three exemplary curves for the cases in which both beams have a radius of 1 mm (blue), 2 mm (red), and 4 mm (green).
-It is evident that the interference visibility can be maintained over a larger lateral shift range in case of larger beam radii.
-In Fig.
-4(b), we show the influence of the wave-vector mismatch on the interference visibility by varying the propagation angle of the second beam relative to the first.
-Here, the angular mismatch has less impact on the interference visibility when both beams have smaller radii, which reflects the general behavior of the last exponential function in Eq.
-(28).
-Figure 4: Interference visibility factor $\eta_{1,2}$ between two light fields.
-(a) Interference visibility resulting from the second beam being laterally shifted relative to the first beam where both beams have a radius of 1 mm (blue), 2 mm (red) or 4 mm (green).
-(b) Dependence of the interference visibility on the relative incidence angle of the second beam [colors correspond to the same beam radii as in (a)].
-(c) Interference visibility of two beams incident on an aperture with varying radius.
-Here, both beams have a radius of 4 mm and one of the beams is laterally shifted by 1 mm from the center of the aperture at normal incidence while the other beam is laterally shifted by 4 mm relative to the first beam and has an incidence angle of 0.01 (fixed parameters).
-In Section LABEL:SISecInterferenceVisibilityFactor of the SI, we discuss quantitatively how interference visibility can be increased by closing an entrance aperture $a$ directly in front of a spectrometer, and we find the conditions for maximal angular mismatch and transverse misalignment between incident beams that still ensure interference.
-In Fig.
-4(c), we display the interference visibility factor in the case where two beams with a radius of 4 mm are incident on an aperture while the size of the aperture is changed.
-One of the beams is laterally shifted by 1 mm relative to the center of the aperture at normal incidence whereas the other beam has an exemplary lateral shift of 4 mm relative to the first beam and is incident at an angle of $0.01^{\circ}$.
-The two beams hit the aperture at a distance of 1 m from their origin, resulting in the second beam being displaced in the plane of the iris by 4.8 mm relative to the center of the iris.
-This configuration leads to a beam overlap region that is displaced relative to the center of the iris, as indicated in the inset of Fig.
-4(c).
-This example shows that closing the aperture (initially considered to have an infinite opening radius) leads to an increase of the interferometric contrast as those parts of the beams that do not overlap get truncated.
-However, when closing the aperture further, the overlapping parts get truncated as well, which leads to reduction in contrast such that there is an optimum aperture opening.
-Note that the slight kink at an aperture radius of 3 mm is due to the transition in how the beam cutting is calculated as discussed in connection with Fig.
-3(b).
 4 Modulations from light–matter interaction
-4.1 First-order non-resonant response
+4.1 First-order response
 The electric-field modifications by linear response of matter is most conveniently implemented by multiplying the incident frequency-domain field, $E_{\text{in}}(\omega)$, with a frequency-domain linear modulation function, $M(\omega)$, to get the output field,
@@
 (13).
-In the case of reflection, finite values for the group-delay dispersion or the third-order phase coefficient may be provided directly to describe, e.g., chirped mirrors.
-4.2 First-order resonant response
-It is convenient to start with a response function in time domain, rather than frequency domain, to treat a resonantly excited material such as an ensemble of molecules in a solvent.
-The frequency-domain modulation function is then obtained by Fourier transformation via Eq.
-(LABEL:SIEqEomegaJFourier) and appropriate further scaling, as described below.
-The linear temporal response function in a Franck–Condon model under the rotating-wave approximation (RWA) is given by [24]
-$S_{\text{RWA}}^{(1)}(t)=\theta(t)\frac{i}{\hbar}\mu_{\text{eg}}^{2}\exp\left(-\gamma t-\frac{1}{2}\Delta\omega^{2}t^{2}\right)e^{-i(\omega_{\text{eg}}-\omega_{0})t}e^{-S_{\text{HR}}}\sum_{j=0}^{N_{\text{vib}}-1}\frac{S_{\text{HR}}^{j}}{j!}e^{-ij\omega_{\text{vib}}t},$ (32)
-[…]
-We now have to introduce an appropriate scaling of $\mu_{\text{eg}}$ to convert between “physical” quantities and computationally stored arrays.
-This problem is similar to the scaling of electric fields that we solved by normalizing with respect to a given temporal power and, ultimately, a given pulse energy, via Eq.
-(LABEL:SIEqTemporalPowerDefinition).
-Thus we apply a similar strategy here.
-Furthermore, we seek a behavior of the modulation function such that longer paths through the material lead to stronger absorption fulfilling the Lambert–Beer law for the pulse energy,
-$W_{\text{out}}=W_{\text{in}}e^{-\alpha L},$ (33)
-with a given extinction coefficient $\alpha$ of the material.
-As a first step, we evaluate Eq.
-(32) on the same time grid as for regular laser pulses [Eq.
-(LABEL:SIEqt(j))] and set $\mu_{\text{eg}}^{2}/\hbar=1$.
-The Fourier transform of the result is the (unscaled) frequency-domain linear susceptibility,
-$\chi_{\text{unscaled}}(\omega)=\operatorname{\mathfrak{F}}\left[\left.S_{\text{RWA}}^{(1)}(t)\right|_{\mu_{\text{eg}}^{2}/\hbar=1}\right],$ (34)
-from which we get the (unscaled) complex index of refraction,
-$\tilde{n}_{\text{unscaled}}(\omega)=\sqrt{\chi_{\text{unscaled}}(\omega)+1},$ (35)
-where it is understood that the square-root symbol indicates a complex-valued square-root operation. Now we obtain a suitable scaling factor,
-$s=\frac{\alpha c}{2\omega_{\alpha}\operatorname{Im}\tilde{n}_{\text{unscaled}}(\omega_{\alpha})},$ (36)
-where $\alpha$ is the absorption coefficient of Eq.
-(33) to which we want to scale the material response and $\omega_{\alpha}$ is the angular frequency at which this absorption coefficient shall be reached (typically a peak of the resulting absorption spectrum that can be found by locating the frequency at which $\operatorname{Im}\tilde{n}_{\text{unscaled}}(\omega)$ has a maximum).
-Using Eq.
-(36) allows us to calculate a scaled susceptibility,
-$\chi(\omega)=s\chi_{\text{unscaled}}(\omega),$ (37)
-and a scaled complex index of refraction,
-$\tilde{n}(\omega)=\sqrt{\chi(\omega)+1},$ (38)
-and finally the desired modulation function,
-$M_{\text{resonant}}(\omega)=\exp\left(i\frac{\tilde{n}(\omega)\omega L}{c}\right),$ (39)
-for a material of thickness $L$.
-[…]
-Figure 5: Linear resonant response. Exemplary modulation function $M_{\text{resonant}}(\omega)$ for an absorbing molecular response (red: real part, blue: imaginary part).
+In the case of reflection, group-delay dispersion or a third-order phase coefficient may be provided directly to describe, e.g., chirped mirrors.
+We treat the case of first-order resonant response, on the example of a molecular sample with an electronic transition coupled to vibrational modes, in Section LABEL:SISecFirstOrderResonantResponse of the SI.
 4.2 Second-order response
@@
 However, we need to consider also the spatial properties of the beams if we want to simulate and understand the effect of beam alignment on nonlinear phenomena.
-Here we derive the explicit equations in terms of the computationally stored complex envelope arrays $\tilde{E}_{t}(j)$ and associated parameters, we take into account the spatial beam profile, and we determine the missing proportionality factor between $E_{\text{s}}(t)$ and $E^{2}(t)$.
-Equation (40) contains multiplicative electric-field terms. Hence, we use Eq. (1) and multiply the propagated fields,
-$\begin{split}&mE_{1,\text{prop}}^{+}(x,y,z,t)E_{2,\text{prop}}^{+}(x,y,z,t)\\ =\,&m\sqrt{\frac{2}{\pi}}\frac{1}{w_{1}(z)}\exp\left[-\frac{(x-x_{1})^{2}+(y-y_{1})^{2}}{w_{1}^{2}(z)}\right]e^{i\mathbf{k}_{1}\cdot\mathbf{r}}\sqrt{\frac{S_{1}}{\delta t}}\tilde{E}_{1,t}(j_{1})e^{-i\omega_{1}(t-T_{1})}\\ &\times\sqrt{\frac{2}{\pi}}\frac{1}{w_{2}(z)}\exp\left[-\frac{(x-x_{2})^{2}+(y-y_{2})^{2}}{w_{2}^{2}(z)}\right]e^{i\mathbf{k}_{2}\cdot\mathbf{r}}\sqrt{\frac{S_{2}}{\delta t}}\tilde{E}_{2,t}(j_{2})e^{-i\omega_{2}(t-T_{2})},\end{split}$ (41)
-wherein the multiplication of the two Gaussian exponentials generates a new Gaussian $A_{\text{p}}(x,y)$ according to Eq.
-(LABEL:SIEqGaussianProduct) of the SI with the product amplitude $A_{\text{p}}$ in Eq.
-(LABEL:SIEqProductAmplitude) and $A_{1}=A_{2}=1$.
-Because we consider the stigmatic approximation, we have $w_{1,x}=w_{1,y}$ and $w_{2,x}=w_{2,y}$ according to Eq.
-(LABEL:SIEqGaussianProduct) to Eq.
-(LABEL:SIEqProductAmplitude) of the SI.
-Further, $m=1$ for second-harmonic generation (SHG) but $m=2$ for sum-frequency generation (SFG) according to Eq.
-(40) due to the two possible permutations $E_{1}E_{2}$ and $E_{2}E_{1}$ that both deliver an SFG field in the same direction.
-We want to write the result of Eq. (41) as a new (but not yet correctly energy-scaled) SFG field,
-$\begin{split}E_{\text{SFG,prop}}^{+}(x,y,z,t)=\,&\sqrt{\frac{2}{\pi}}\frac{1}{w_{\text{SFG}}}\exp\left[-\frac{(x-x_{\text{SFG}})^{2}+(y-y_{\text{SFG}})^{2}}{w_{\text{SFG}}^{2}}\right]e^{i\mathbf{k}_{\text{SFG}}\cdot\mathbf{r}}\\ &\times\sqrt{\frac{S_{\text{SFG,unscaled}}}{\delta t}}\tilde{E}_{\text{SFG},t}(j_{\text{SFG}})e^{-i\omega_{\text{SFG}}(t-T_{\text{SFG}})},\end{split}$ (42)
-with appropriate parameters. Comparing the individual factors of Eq. (41) and Eq. (42), we find:
-[…]
-[…]
-[…]
-$\omega_{\text{SFG}}$ $=\omega_{1}+\omega_{2},$ (46)
-$\mathbf{k}_{\text{SFG}}$ $=\mathbf{k}_{1}+\mathbf{k}_{2},$ (47)
-$T_{\text{SFG}}$ $=\frac{\omega_{1}T_{1}+\omega_{2}T_{2}}{\omega_{1}+\omega_{2}},$ (48)
-$\tilde{E}_{\text{SFG},t}(j_{\text{SFG}})$ $=\tilde{E}_{1,t}(j_{1})\tilde{E}_{2,t}(j_{2}),$ (49)
-$\sqrt{\frac{2}{\pi}}\frac{1}{w_{\text{SFG}}}\sqrt{\frac{S_{\text{SFG,unscaled}}}{\delta t}}$ $=m\frac{2}{\pi}\frac{1}{w_{1}(z)w_{2}(z)}A_{\text{p}}\frac{\sqrt{S_{1}S_{2}}}{\delta t}.$ (50)
-Note that the beam radius of the SFG beam is smaller than those of the fundamental beams, as expected due to the field multiplication.
-If both incident beams have the same radius, they are larger than the SFG radius by a factor of $\sqrt{2}$ according to Eq.
-(45).
-Solving Eq.
-(50) for $S_{\text{SFG,unscaled}}$, using the explicit results for $w_{\text{SFG}}$ from Eq.
-(45) and for $A_{\text{p}}$ from Eq.
-(LABEL:SIEqProductAmplitude) (with $A_{1}=A_{2}=1$) and simplifying leads to
-$S_{\text{SFG,unscaled}}=\frac{2m^{2}}{\pi}\frac{1}{w_{1}^{2}(z)+w_{2}^{2}(z)}\frac{S_{1}S_{2}}{\delta t}\exp\left[-2\frac{(x_{1}-x_{2})^{2}+(y_{1}-y_{2})^{2}}{w_{1}^{2}(z)+w_{2}^{2}(z)}\right].$ (51)
-We illustrate how $S_{\text{SFG,unscaled}}$ varies when displacing one of the fundamental beams relative to the other in Fig.
-6.
-Each of the fundamental beams has a pulse energy of 1 mJ and the beam radius of one beam was set to 1 mm (blue), 2 mm (red), and 4 mm (green) while that of the other beam is kept at 1 mm.
-It is evident that the scaling factor generally shows the largest change with increasing overlap when both beams have small radii.
-In other words, the SFG intensity at maximum overlap increases with increasing energy density of the fundamental beams.
-Figure 6: Beam energy scaling factor for sum-frequency generation.
-The plots show the dependence of the beam energy scaling factor for sum-frequency generation, $S_{\text{SFG,unscaled}}$, on the lateral shift between two fundamental beams.
-The beam radius of the first beam was varied (blue: 1 mm, red: 2 mm, green: 4 mm), whereas that of the second beam was held constant (1 mm).
-Concerning the multiplication of envelopes in Eq.
-(49), we have to use the appropriate array indices.
-Note that we have to add the respective propagation times $T_{1}$ and $T_{2}$ to Eq.
-(LABEL:SIEqt(j)) before extracting the relevant indices $j_{1}$ and $j_{2}$ of the array elements to be multiplied in Eq.
-(49).
-In the multiplication, it will thus happen that a theoretically requested array index is out of range of the available indices $j=\{0,1,\ldots,N_{\text{s}}-1\}$, i.e., the time shift between the two pulses is such that for a time coordinate $t(j_{1})$ at index $j_{1}$ of the first field no “matching” index $j_{2}$ is found because the particular $t(j_{1})$ is not contained in the list for $t(j_{2})$ of the second field.
-Then we assume the multiplier field to be zero, and thus also the SHG field will vanish at this time coordinate.
-Now we have to find an appropriate scaling factor for the SFG field.
-In principle, proper perturbation theory of light–matter interaction delivers already the correct scaling factor of the absolute signal strength in terms of an appropriate response function.
-This is, however, not applicable in our situation because we chose a simplified treatment for reasons of computational speed, rather than dealing with a more realistic but much more complicated full spatial–temporal model [40, 41, 42, 43, 44].
-In particular, the response-function treatment provides the field at one particular spatial position only, and one then has to solve the wave equation with that nonlinear field as a source term upon propagation through the full length of the material.
-This is beyond our scope.
-Thus, we do not take into account propagation effects and want the final signal to scale simply with the square of the incident field.
-Then we have to ensure, however, that the pulse energy in the second-harmonic field does not exceed the sum of pulse energies in the initial fields because then energy conservation could not be fulfilled.
-In practice, we will require only a (small) fraction of the power to be converted in order to prevent saturation effects.
-We derive in Section LABEL:SISecAbsoluteAmplitudeScalingOfSecondOrderFields of the SI an appropriate scaling factor.
-If we allow only a fraction $\eta_{2}$ of the fundamental pulse energy $W_{0}$ to be converted to SHG under optimum conditions (leading to a maximum SHG pulse energy $W_{\text{SHG,max}}$), the final result for the pulse-energy scaling factor is
-$S_{\text{SFG}}=\eta_{2}\frac{W_{0}}{W_{\text{SHG,max}}}S_{\text{SFG,unscaled}}.$ (52)
-We may set $0<\eta_{2}\leq 1$ at will to guarantee energy conservation in conjunction with an appropriate reduction of the fundamental energies as explained in Section 4.5.
-We may select, e.g., $\eta_{2}=10\,\mathrm{\%}$ to stay away from the saturation regime.
-This evaluation has to be done only once for a given laser (unless the laser output parameters are changed).
-The intermediate multiplication result $\eta_{2}W_{0}/W_{\text{SHG,max}}$ can then simply be stored as a number and used as a multiplier on $S_{\text{SHG,unscaled}}$.
-4.4 Nonlinear signal beam curvature
-We now discuss the beam radius and radius of curvature of a nonlinear signal beam that arises from the corresponding properties of the incident fields.
-The incident fields may be mutually different in general, e.g., in pump–probe experiments with different beam radii of pump and probe beams.
-We thus need to address where the focus of the resulting beam will be located.
-A detailed derivation is presented in Section LABEL:SISecDerivationOfNonlinearSignalBeamCurvature of the SI, and we only quote the results here.
-It turns out that the $n$th-order nonlinear signal beam radius, $w_{\text{s}}(z)$, at the position of a thin sample, $z$, is obtained from the individual beam radii of the $n$ incident beams, $w_{i}(z)$, via
-$\frac{1}{w_{\text{s}}^{2}(z)}=\sum_{i=1}^{n}\frac{1}{w_{i}^{2}(z)},$ (53)
-in generalization of Eq. (45). The curvature radius of the signal beam at the sample position, $R_{\text{s}}(z)$, is given by
-$R_{\text{s}}(z)=\frac{\pi w_{\text{s}}^{2}(z)}{\lambda\zeta_{\text{s}}}$ (54)
-with the longitudinal position parameter
-$\zeta_{\text{s}}=w_{\text{s}}^{2}(z)\sum_{i=1}^{n}\frac{\alpha_{i}\zeta_{i}}{w_{i}^{2}(z)}$ (55)
-with phase-matching coefficients $\alpha_{i}\in\{-1,+1\}$ of the incident beams indicating with which sign they contribute to the calculation of the nonlinear signal phase-matching direction, and
-$\zeta_{i}=-\frac{\operatorname{Re}q_{i}}{\operatorname{Im}q_{i}}$ (56)
-using the complex beam parameters $q_{i}$ of the incident beams [compare Eq. (LABEL:SIEqComplexRadiusOfCurvature)].
-Figure 7: Nonlinear signal beam.
-(a) Lateral position of the SHG beam waist emerging from two collinear fundamental beams with the same radius while the second beam is laterally shifted.
-The lateral position of the SHG beam is given relative to that of the first fundamental beam.
-(b) Beam waist of the SHG beam depending on the beam waist of the second fundamental beam.
-The waist of the first fundamental beam is set to 1 mm.
-For one particular limiting case, consider that the sample or nonlinear crystal is placed near the position of the beam waist of all incident beams, i.e., $\zeta_{i}\ll 1$.
-This represents the “plane-wave” situation (for which $R_{i}\to\infty$) that can be realized either by placing the sample at the focal point or by employing collimated beams.
-Then, $\zeta_{\text{s}}\ll 1$, and the signal focus is located at the same longitudinal position as the incident beams are.
-In other words, the signal beam is created at its waist in the sample.
-In Fig.
-7(a), we illustrate different focusing conditions of two incoming beams by plotting the lateral SHG beam waist position while varying the lateral displacement between the incoming beams.
-We consider two collinear beams under perfect phase matching.
-In case of perfect co-focusing, the waist position of the SHG beam is the same as that of the incident beams.
-When one of the incident beams is laterally displaced, the SHG beam waist position also shifts, but by half the value of the displacement.
-For SHG, the Rayleigh range $z_{\text{R}}$ is the same as that of the fundamental [22].
-Thus, the half-divergence angle of SHG is smaller by $1/\sqrt{2}$ with respect to the fundamental according to Eq.
-(LABEL:SIEqHalfDivergenceAngle), with the beam waist smaller by the same factor.
-If the two beams are co-focused on the sample while the first beam has a fixed radius and the radius of the other beam is increased, the radius of the SHG beam approaches that of the first beam asymptotically [Fig.
-7(b)].
-Consider for further illustration, as another exemplary special case, pump–probe transient absorption spectroscopy, even though its response-function treatment is beyond the scope of the present work.
-Nevertheless, this case is instructive for illuminating the generality of the derived beam-parameter results.
-In that case, the phase-matching direction of the signal is given by
-$\mathbf{k}_{\text{s}}=-\mathbf{k}_{\text{pu}}+\mathbf{k}_{\text{pu}}+\mathbf{k}_{\text{pr}}$ (57)
-for incident pump ($\mathbf{k}_{\text{pu}}$) and probe ($\mathbf{k}_{\text{pr}}$) wave vectors.
-The wave vector $\mathbf{k}_{\text{pu}}$ appears twice in this equation, once with a positive sign and once with a negative sign, because there are two (phase-conjugate) interactions with the pump pulse.
-Thus, the phase-matching coefficients are $\{\alpha_{1},\alpha_{2},\alpha_{3}\}=\{-1,+1,+1\}$ and the signal propagates into the direction of the probe beam (leading to constructive or destructive interference that is then interpreted in measurements as absorbance changes).
-From Eq.
-(53), we obtain
-$\frac{1}{w_{\text{s}}^{2}(z)}=\frac{2}{w_{\text{pu}}^{2}(z)}+\frac{1}{w_{\text{pr}}^{2}(z)}.$ (58)
-It is often customary to choose the pump-beam radius (much) larger than the probe-beam radius.
-Under this condition, we find from Eq.
-(58) that $w_{\text{s}}\approx w_{\text{pr}}$ and from Eq.
-(55) that $\zeta_{\text{s}}\approx\zeta_{\text{pr}}$, and thus the signal beam has the exact same characteristics as the probe beam.
-This is the reason for choosing a larger pump-beam diameter because then the interference leads to a spatially uniform transient absorption signature across the full probe-beam profile.
-Otherwise, the signal beam might deviate from the probe beam and one loses contrast.
-4.5 Reduction of fundamental pulse energies upon nonlinear signal generation
+In Section LABEL:SISecDerivationSecondOrderResponse of the SI, we derive the explicit equations in terms of the computationally stored complex envelope arrays $\tilde{E}_{t}(j)$ and associated parameters, we take into account the spatial beam profile, and we determine the missing proportionality factor between $E_{\text{s}}(t)$ and $E^{2}(t)$.
+The correct description of the nonlinear signal field further requires to calculate its beam radius and the radius of curvature, depending on the incident fields.
+These relations are derived in Section LABEL:SISecNonlinearSignalBeamCurvature of the SI.
 When new signal fields are generated according to second-order or third-order response and no energy is absorbed in the nonlinear crystal or sample, the total energy summed up over all laser beams should be conserved.
@@
 Otherwise, it might happen that the total energy of all output beams after a nonlinear GOE exceeds the total energy of all incident beams.
-We arrive at the correct energy, contributed to the new beam by each fundamental, by analyzing the relevant number of photons for each microscopic signal-generation process.
-Let us consider sum-frequency generation (SFG) as in Eq.
-(40).
-Each SFG photon has an energy of $\hbar\omega_{\text{SFG}}$, and the fundamental beams $\tilde{E}_{1}(t)$ and $\tilde{E}_{2}(t)$ carry photons of energy $\hbar\omega_{1}$ and $\hbar\omega_{2}$, respectively.
-The incident fundamental pulse energies, $W_{1,\text{in}}$ and $W_{2,\text{in}}$, and the generated SFG pulse energy, $W_{\text{SFG}}$, can be obtained from the fields via Eq.
-(LABEL:SIEqEnergyIsScaleTimesNormSquared) from the SI.
-Thus, a number of $W_{\text{SFG}}/(\hbar\omega_{\text{SFG}})$ photons have to be generated.
-This requires the same number of photons contributed from each of the incident beams, reducing their energies accordingly by
-$\Delta W_{1}=\frac{\hbar\omega_{1}}{\hbar\omega_{\text{SFG}}}W_{\text{SFG}},\quad\Delta W_{2}=\frac{\hbar\omega_{2}}{\hbar\omega_{\text{SFG}}}W_{\text{SFG}}.$ (59)
-Analogously, one obtains the beam reductions for second-harmonic generation by setting $W_{2}=W_{1}$, i.e., the fundamental beam is reduced by twice the value (in addition to the reduction from potential SFG contributions).
-Furthermore, it is possible that more than two beams interact in any given GOE with second-order response.
-In that case, we calculate the reductions for all beam-pair combinations (SFG and SHG) and then, in the end, obtain the reduced fundamental output pulse energies by
-$W_{1,\text{out}}=W_{1,\text{in}}-\sum_{k}\Delta W_{1,k},\quad W_{2,\text{out}}=W_{2,\text{in}}-\sum_{k}\Delta W_{2,k},$ (60)
-where the index $k$ runs over all pairwise combined processes.
-Note that we reduce the fundamental pulse energies only after having generated all second-order beams with the full incident energy.
-This ensures that no particular pair of beams is “prioritized” and has a higher pulse energy than any of the others when calculating the SFG and SHG beams successively.
-Of course the beam-energy reductions occur in addition to those from beam clipping as discussed in Section 3.1, and we do not modify $W$ directly but rather the pulse-energy scaling factor $S$ from Eq.
-(LABEL:SIEqEnergyIsScaleTimesNormSquared).
-Working with photon numbers and center frequencies only is an approximation.
-In principle, one could analyze the complete spectral intensity shape for the correct result [32], but since we require only the adjustment of absolute energies here, such small variations are irrelevant.
-For the spectral–temporal shapes of the signal and transmitted pulses themselves we take into account the explicit field profiles.
+We arrive at the correct energy, contributed to the new beam by each fundamental, by analyzing the relevant number of photons for each microscopic signal-generation process as derived in Section LABEL:SISecReductionOfFundamentalPulseEnergiesUponNonlinearSignalGeneration of the SI.
 5 Graphical laser beam representation
-Simulating linear and nonlinear optical phenomena leads to laser beams propagating through space with parameters as determined in Section 2, Section 3, and Section 4.
+Simulating linear and nonlinear optical phenomena leads to laser beams propagating through space with parameters as determined in Section 2 and Section 4.
 If one only seeks numerical results, then the calculation is already complete because the obtained parameters provide a full characterization.
@@
 Instead, we discuss some general points that are relevant for any such three-dimensional modeling that is sufficiently fast and realistic.
-5.1 Gaussian beam scaling
-In this section, we investigate general spatial properties of Gaussian beams.
-The question will be how the shapes of Gaussian laser beams are related to each other given arbitrary beam parameters.
-We will find that any Gaussian beam can be derived from the general shape of a “reference beam” by suitable scaling.
-Thus, it is not necessary to re-calculate, point by point, the spatial envelopes of a beam for any new set of beam parameters, but one can simply generate a Gaussian model once, as will be shown in Section 5.2, and then scale it accordingly.
-Let us assume that we start with a suitable representation of a Gaussian “reference beam” that is given by the beam radius,
-$\tilde{w}(z)=\tilde{w}_{0}\sqrt{1+\left(\frac{z}{\tilde{z}_{\text{R}}}\right)^{2}},$ (61)
-as a function of the distance $z$ from the beam waist with a “Rayleigh length”
-$\tilde{z}_{\text{R}}=\frac{\pi\tilde{w}_{0}^{2}}{\lambda M^{2}},$ (62)
-for propagation along the $\hat{\mathbf{z}}$ direction, wavelength $\lambda$, and beam-quality parameter $M^{2}\geq 1$ (for Gaussian beams, $M^{2}=1$), with the waist $\tilde{w}_{0}$ located at $z=0$.
-The tilde symbolizes the “reference” character.
-We seek a new beam
-$w(z)=w_{0}\sqrt{1+\left(\frac{z-z_{0}}{z_{\text{R}}}\right)^{2}}$ (63)
-with a new Rayleigh length
-$z_{\text{R}}=\frac{\pi w_{0}^{2}}{\lambda M^{2}}$ (64)
-whose (potentially different) waist $w_{0}$ is located at position $z=z_{0}$, and we would like to obtain this beam as a suitably transformed version of the reference beam.
-It is often convenient to treat Gaussian beam propagation using a complex-valued beam curvature radius $q$, as introduced in Eq. (LABEL:SIEqComplexRadiusOfCurvature) in the SI, that is defined as [45]
-$\frac{1}{q(z)}=\frac{1}{R(z)}+i\frac{\lambda M^{2}}{\pi w^{2}(z)}$ (65)
-at a position $z$ with relation to the wavefront curvature radius $R$. Beam propagation in free space follows [46, 47]
-$q(z)=q_{0}+z$ (66)
-for propagation by $z$, and $q_{0}=q(0)$ [see also Eq. (LABEL:SIEqPropagationFreeSpace) in the SI]. Optical elements such as lenses or mirrors will lead to a modification of $q$ [14].
-According to Eq. (66), we have
-$\tilde{q}(z)=\tilde{q}_{0}+z$ (67)
-and
-$\tilde{q}_{0}=-i\tilde{z}_{\text{R}}$ (68)
-for the propagation law of the reference beam, and its complex curvature at its waist position, respectively.
-Assume we are given $q=\operatorname{Re}q+i\operatorname{Im}q$, with $\operatorname{Re}q$ and $\operatorname{Im}q$ providing the real and imaginary parts of the complex radius of curvature $q$ of the “new” beam, immediately after transfer through an optical element.
-Then the new beam waist is located at a position $z_{0}=-\operatorname{Re}q$, translated with respect to the current position at the optical element, and the transformed Rayleigh length is given by $z_{R}=-\operatorname{Im}q$.
-Let us define the scaling factor
-$b=\frac{z_{\text{R}}}{\tilde{z}_{\text{R}}}=-\frac{\operatorname{Im}q}{\tilde{z}_{\text{R}}}$ (69)
-such that
-$z_{\text{R}}=b\tilde{z}_{\text{R}}$ (70)
-and
-$\frac{w_{0}}{\tilde{w}_{0}}=\frac{z_{\text{R}}^{2}}{\tilde{z}_{\text{R}}^{2}}=b^{2}.$ (71)
-Then we obtain
-$w(z+z_{0})$ $=w_{0}\sqrt{1+\left(\frac{z}{z_{\text{R}}}\right)^{2}}$ (72)
-$=b^{2}\tilde{w}_{0}\sqrt{1+\left(\frac{z}{b\tilde{z}_{\text{R}}}\right)^{2}}$ (73)
-$=b^{2}\tilde{w}\left(\frac{z}{b}\right).$ (74)
-Thus, the new beam $w$ can be obtained from the reference beam $\tilde{w}$ by a scaling factor $b^{2}$ along the transverse coordinates ($x$ and $y$) and by a scaling factor $b$ along the longitudinal propagation coordinate $z$ [noting that the beam is increased in size for $b>1$ along transverse as well as longitudinal coordinates despite the factor $b$ appearing once in the numerator and once in the denominator of Eq.
-(74)].
-The result is then shifted by $-z_{0}$, i.e., by $+\operatorname{Re}q$, along the positive propagation direction.
+In Section LABEL:SISecGaussianBeamScaling of the SI, we show that any Gaussian beam can be derived from the general shape of a “reference beam” by suitable scaling.
+Thus, it is not necessary to re-calculate, point by point, the spatial envelopes of a beam for any new set of beam parameters, but one can simply generate a Gaussian model once, and then scale it accordingly.
+Let us assume we have available a representation of a Gaussian “reference beam” that is given by its beam radius $\tilde{w}(z)$ at position $z$ along the propagation direction.
+From that reference, we can obtain the representation of any new Gaussian beam, $w(z)$, according to Eq.
+(LABEL:SIEqMeshTransformation),
+$w(z+z_{0})=b^{2}\tilde{w}\left(\frac{z}{b}\right),$ (15)
+by applying a scaling factor $b^{2}$ along the transverse coordinates ($x$ and $y$) and a scaling factor $b$ along the longitudinal propagation coordinate $z$.
+The result is then shifted by $-z_{0}$, along the positive propagation direction, to obtain the correct beam-waist position.
+The scaling factor
+$b=\frac{z_{\text{R}}}{\tilde{z}_{\text{R}}}$ (16)
+is the ratio of the Rayleigh lengths of new beam and reference beam [Eq.
+(LABEL:SIEqScalingFactorRayleighLengths)].
 For representing finite-length beams, one can “clip” the resulting object appropriately at its end positions such that only the relevant section (e.g., between two mirrors) is displayed.
-5.2 Three-dimensional beam model
-Now that we have seen in Section 5.1 how any Gaussian laser beam can be obtained from one given reference shape, we discuss how such a reference shape can be prepared efficiently.
-In three-dimensional computer graphics, solid objects are generally modeled as polygon meshes, i.e., a set of points in three-dimensional space (called “vertices”) that are connected by straight lines (“edges”) that in turn form small planes (“faces”).
-The collection of vertices and lines is sought such that the resulting set of faces approximates the real surface of the given object.
+We now discuss how such a reference shape can be prepared efficiently.
+In three-dimensional computer graphics, solid objects are generally modeled as polygon meshes, i.e., a set of points (“vertices”) in three-dimensional space that are connected by straight lines (“edges”) that in turn form small planes (“faces”).
+The collection of vertices and edges is sought such that the resulting set of faces approximates the real surface of the given object.
 In practice, striving for good performance in real-time rendering applications, one seeks to keep the number of vertices as small as possible to reduce calculation time, yet still capturing the essential shape of the underlying object as accurately as possible.
-In the present case, we desire to find the minimum number of vertices and their positions that can be used to represent a Gaussian laser beam.
-The challenge will be that the beam radius of a Gaussian evolves as a hyperbola along the propagation direction as given by Eq.
-(61) and as a circle around the circumference with radius $w$, while the edges of the mesh are straight line segments.
-For optimizing the number and distribution of vertices, we introduce a local relative length error, $\varepsilon$, defined as the relative length difference between a curve segment following the real Gaussian shape and a straight edge, connecting the same two neighboring vertices.
-Then we demand this relative error to be below a certain threshold for all the edges in the mesh, and we seek a vertex distribution that minimizes the number of required vertices for a user-specified error.
-We derive in Section LABEL:SISecOptimalVertexDistribution of the SI the minimum number of vertices along the propagation direction for any Gaussian beam, independent of its beam parameters, and find that just five vertices along the propagation direction are sufficient to ascertain a relative maximum error of $\varepsilon=1\,\mathrm{\%}$.
-This signifies that in many cases, Gaussian beams are well approximated by straight lines at long distances, i.e., the asymptotes of the hyperbolic divergence that is found for $z\gg\tilde{z}_{\text{R}}$ in Eq.
-(61), and that for the curving sections (between the Rayleigh points) few additional vertices are sufficient.
-Around the circumference, we find that $13$ vertices are required for $\varepsilon=1\,\mathrm{\%}$, signifying that, compared to the longitudinal direction, more straight edges are required to approximate a circle.
-Altogether, thus, one needs a minimum of $5\times 13=65$ vertices.
-Requiring an error of $\varepsilon=0.5\,\mathrm{\%}$ changes these numbers to $7\times 18=126$.
-In practice, it may not be practical to use so few vertices along the longitudinal direction because then the resulting mesh might be much larger than the required beam segment between two optical elements.
-In that case, there would be no mesh vertices at all in the required spatial region.
-While it is still possible, using suitable rendering agorithms with clipping, to display objects whose vertices are all located outside of the visible region, it may become preferable to define the object with closer-spaced vertices.
-In particular, the mesh vertices may be required not only for a faithful representation of an (infinitely long) Gaussian beam, but they are also used to terminate the graphical representation when the laser intersects with an object.
-For that reason, it is helpful to adopt a smaller-step longitudinal mesh spacing in practice by selecting an appropriate maximum step size between any two vertices along the propagation direction, i.e., by selecting an appropriate $\Delta z_{\text{max}}$ in Eq.
-(LABEL:SIEqOptimalLongitudinalMeshSpacing).
-The present analysis assures, however, that the relative local error $\varepsilon$ will always be smaller than the given error threshold, and thus the graphical output will approximate the real beam shape well.
-5.3 Laser color
+In Section LABEL:SISecOptimalVertexDistribution of the SI, we find the minimum number of vertices and their positions that can be used to represent a Gaussian laser beam for a given maximum allowed error.
+That error is defined as the relative length difference between a curve segment following the real Gaussian shape and a straight edge, connecting the same two neighboring vertices.
+In particular, we found that the spatial envelope of any Gaussian beam can be visualized, at a $1\,\mathrm{\%}$ error level, in three-dimensional space using a polygonal mesh with just five vertices connected by straight edges along the propagation direction to describe the hyperbolic beam evolution and $13$ vertices along the circumference to approximate the circular symmetry.
+Lastly, let us discuss the perceived color of a laser beam that is relevant in realistic graphical representations.
 When a laser pulse is scattered off a surface, an observer can perceive the spot with a certain color and a certain luminance (“brightness”) that depends on the pulse spectrum, energy, beam radius, and surface properties.
-For a realistic visualization of such a cross section or of the full laser-beam mesh discussed in Section 5.1 and Section 5.2, we should assign a corresponding computer-graphics color code.
-For this purpose, we adapt procedures from computer graphics visualization.
+For a realistic visualization of such a cross section or of the full laser-beam mesh discussed in Sections LABEL:SISecGaussianBeamScaling and LABEL:SISecOptimalVertexDistribution of the SI, we assign a corresponding computer-graphics color code.
 In Section LABEL:SISecColorPerception of the SI, we use a “standard observer” model of average human color perception as defined by the “Commission Internationale de l’Éclairage” (CIE).
-We use the XYZ standard in its most recent implementation [48] and employ the “2-deg XYZ color matching functions transformed from the CIE (2006) 2-deg LMS cone fundamentals” to get the XYZ color values for the given laser spectrum.
+We use the XYZ standard in its most recent implementation [36] to get the XYZ color values for the given laser spectrum.
 This color code is subsequently transformed to an RGB representation that can be displayed in computer graphics.
@@
 Having introduced our model for linear and nonlinear optics simulations, we now show exemplary simulation results.
-For that purpose, we consider a setup for a well-known pulse-characterization method, SHG frequency-resolved optical gating (FROG) [49, 26].
+For that purpose, we consider a setup for a well-known pulse-characterization method, second-harmonic generation (SHG) frequency-resolved optical gating (FROG) [37, 26].
 The setup as shown in Fig.
-8(a) is based on a Mach–Zehnder-type interferometer.
+1(a) is based on a Mach–Zehnder-type interferometer.
 Note that, to achieve perfect interferometer alignment, the GOEs shown in Fig.
-8(a) were placed programmatically by setting numerical position values rather than by placing them “by hand” in VR.
+1(a) were placed programmatically by setting numerical position values rather than by placing them “by hand” in VR.
 The laser emitter in the top-left corner of Fig.
-8(a) generates Gaussian pulses with a center wavelength of 800 nm and $\tau_{\text{p}}=15$ fs duration [intensity full width at half maximum as in Eq.
+1(a) generates Gaussian pulses with a center wavelength of 800 nm and $\tau_{\text{p}}=15$ fs duration [intensity full width at half maximum as in Eq.
 (LABEL:SIEqNonNormalizedTemporalEnvelope)].
@@
 Under the assumption of perfect phase matching, the second-order response of the BBO was modeled to be “complete,” i.e., the second harmonics of the two individual beams in $2\mathbf{k}_{1}$ and $2\mathbf{k}_{2}$ direction are generated in addition to the “collaborative” SHG in $\mathbf{k}_{1}+\mathbf{k}_{2}$ direction [Fig.
-8(b)].
+1(b)].
 The latter was detected by a spectrometer.
-In the scan procedure in VR, the delay stage was incremented from $-100$ fs to $+100$ fs in steps of $0.5$ fs.
+In the scan procedure in VR, the delay stage was incremented from $-100$ fs to $+100$ fs in steps of $0.2$ fs.
 Rather than directly calculating the resulting FROG trace from an idealized “FROG signal equation,” the data was acquired analogously to reality, i.e., after moving the delay stage to a new step position, the SHG spectrum resulting from the interference of $\mathbf{k}_{1}$ and $\mathbf{k}_{2}$ in the BBO was calculated according to Section 4.2 before moving to the next step position.
-The entire scan took $22\,\mathrm{s}$, and the consistently high and stable frame rate allowed real-time monitoring of changes in the fringe pattern detected by the spectrometer at each scanning step.
+The entire scan took $45\,\mathrm{s}$, and the consistently high and stable frame rate allowed real-time monitoring of changes in the fringe pattern detected by the spectrometer at each scanning step.
 The thus-obtained “experimental” FROG trace shown in Fig.
-8(c) was evaluated using a commercial FROG analysis program [50], yielding the reconstruction results shown in Fig.
-8(d–f).
+1(c) was evaluated using a commercial FROG analysis program [38], yielding the reconstruction results shown in Fig.
+1(d–f).
 Using a grid size of 64, the minimum FROG error was $2.5\times 10^{-4}$.
 Since simulated data have been used, the reconstruction quality is excellent [compare Fig.
-8(c) and Fig.
-8(d)].
+1(c) and Fig.
+1(d)].
 For a more immersive experience, artificial noise could be added to the acquired data.
@@
 In the present work, we have described the development of a model for real-time simulation of linear and nonlinear optical phenomena and spectroscopy.
-The term “real time” means that we endeavor to find such a level of approximation that allows calculation times of complete experimental configurations on the order of $10\,\mathrm{ms}$ on consumer-grade hardware.
+The term “real time” means that we endeavor to find such a level of approximation that allows calculation times of complete experimental configurations on the order of $8\,\mathrm{ms}$ on consumer-grade hardware.
 We chose to implement a mixture of geometrical optics and Gaussian wave optics because the former offers computational speed and simplicity, while the latter provides the correct evolution of beam radius and beam curvature during propagation.
 A finite beam radius is in turn important to describe effects of beam overlap in interference phenomena and nonlinear signal generation.
-A summary of features and limitations is given in tabular form in Section LABEL:SISecFeaturesAndLimitations of the SI.
-In particular, we provided expressions that can be used to computationally process frequency- and time-sampled electric field evolutions of femtosecond laser pulses.
+We summarize in Table 1 basic features and consequences of the physical model of the present work. This list is not meant to be exhaustive but rather to capture essential points for a quick overview.
+Table 1: Features and limitations of physical model.
+Property Choice Consequence
+Beam profile Gaussian Finite beam overlap included
+Beam quality $M^{2}$ factor Real beam divergence approximated
+Beam symmetry Radially symmetric No astigmatism
+Beam clipping No diffraction Geometrical effects included; no gratings
+Beam focusing Ideal No chromatic or spherical aberration
+Polarization direction Scalar electric fields No birefringence; no polarization
+Electric field Discrete sampling Spatial–temporal interference included
+First-order response Response function Dispersion and absorption included
+Second-order response Instantaneous Pulse-shape effects included
+We provided expressions that can be used to computationally process frequency- and time-sampled electric field evolutions of femtosecond laser pulses.
 We removed fast phase oscillations due to the carrier frequency (from the time-domain field) and due to propagation (from the frequency-domain field) to facilitate numerical stability.
-In addition, a suitable normalization condition separates the pulse energy and the amplitude changes that occur due to beam-radius evolution.
+In addition, we separated the pulse energy from the amplitude changes that occur due to beam-radius evolution.
 We derived equations that can be used to calculate geometric beam parameters after transmission of laser beams through finite circular apertures, ignoring astigmatism and diffraction.
@@
 Laser pulses are modified by passing through GOEs.
-In particular, we considered the response-function formalism to describe linear non-resonant dispersion, resonant absorption (using a Franck–Condon model), and second-order non-resonant response.
+We used the response-function formalism to describe linear non-resonant dispersion, resonant absorption (using a Franck–Condon model), and second-order non-resonant response.
 The latter allowed inclusion of sum-frequency and second-harmonic generation.
-While such processes had been treated extensively in the scientific literature, we derived simplified approximative results that nevertheless take into account the Gaussian beam overlap between generating pulses and the Gaussian beam curvatures (i.e., focusing properties), and how both properties translate to the nonlinearly generated signal beam.
-We derived general spatial properties of Gaussian laser beams that are relevant for graphical representation.
-In particular, we found that the spatial envelope of any Gaussian beam can be visualized, at a $1\,\mathrm{\%}$ error level, in three-dimensional space using a polygonal mesh with just five vertices connected by straight edges along the propagation direction to describe the hyperbolic beam evolution and 13 vertices along the circumference to approximate the circular symmetry.
-Such a rendered object can then be scaled to represent Gaussian beams with any other spatial parameters, and we derived the scaling factors.
-For color visualization, we also obtained the relation between the physical laser spectrum and the computer-graphics RGB color system.
+While such processes had been treated extensively in the literature, we derived simplified approximate results that nevertheless take into account the Gaussian beam overlap between generating pulses and the Gaussian beam curvatures (i.e., focusing properties), and how both properties translate to the nonlinearly generated signal beam.
+We derived spatial properties of Gaussian laser beams for graphical representation.
+We found the minimum number of vertices of a polygon mesh that keeps the local relative error below a given threshold.
+Such an object can be scaled to represent Gaussian beams with any other spatial parameters, and we derived the scaling factors.
+For color visualization, we also obtained the relation between the physical laser spectrum and the computer-graphics RGB(A) color system.
 The model described in the current work forms the core of an interactive and immersive virtual-reality (VR) simulation of an ultrafast laser laboratory that we have recently developed [14, 15].
@@
 […]
+Author contributions T.B.
+derived the equations, coded the physical model, and wrote the paper.
+S.M.
+coded didactic missions, carried out and analyzed the simulations, and prepared the figures.
+A.M.
+coded the femtoPro core software and framework.
+T.B.
+and S.v.M.
+supervised the project.
+All authors discussed the results and edited the manuscript.
+[…]
 Acknowledgment We thank Pavel Malý for deriving equations of Gaussian field overlap and nonlinear signal-beam curvature.
@@
 […]
-Data Availability Statement The data that support the findings of this study are openly available at the following DOI: [link will be inserted upon publication].
+Data Availability Statement The data that support the findings of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.15535423 [39].
 […]